### v1

# Scalable LLM Math Reasoning Acceleration with Low-rank Distillation

Abstract
Due to long generations, large language model (LLM) math reasoning demands significant computational resources and time. While many existing efficient inference methods have been developed with excellent performance preservation on language tasks, they often severely degrade math performance. In this paper, we propose Caprese, a resource-efficient distillation method to recover lost capabilities from deploying efficient inference methods, focused primarily in feedforward blocks. With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the math capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs. Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>11% reduction to generate 2048 tokens with Qwen 2.5 14B) while encouraging response brevity.
1 Introduction
With the increasing capabilities of large language models (LLMs) [VSP+17], evaluations are also becoming increasingly sophisticated, typically involving multi-step reasoning such as in math problem solving. These tasks tend to demand long generation which drives up latency, making efficiency a dire issue. Fortunately, many efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization. However, for math reasoning tasks, many efficient inference algorithms begin to break down, decimating performance, despite their robustness in language settings. Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and math capabilities.
One of the main differences between math reasoning and many language tasks is the generation length. Reasoning typically involves long generation due to improved performance from chain-of-thought (CoT) [WWS+22], which often surpasses the length of the input query. However, CoT performance falls apart with efficient algorithms that introduce approximation errors which build up over time. For instance, deploying CATS [LLZ+24], a sparse thresholding method, on Gemma 2 2B [TRP+24] has little impact on language generation performance, but knocks GSM8K [CKB+21] accuracy from 51.02% to 34.42% (Table 1). Similarly, for thinking models: applying GRIFFIN [DCC24], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [GYZ+25] drops MATH-500 [LKB+23] accuracy from 79.40% to 42.00% (Table 2). Moreover, since generation is much more demanding computationally than prefill (e.g., >150$\times$ slower to generate than prefill 1K tokens on one GPU with Llama 3.1 8B), there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling. Thus, this poses two key inference challenges with math reasoning:
  1. 1.
    Instability: Long CoTs can be sensitive to mistakes. Even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [ZCX+24]. Hence, efficient inference algorithms should respect the delicacy of CoT.
  2. 2.
    Inefficiency: Due to the reliance on CoT, which tend to lead to long generation, the already computationally expensive LLM autoregressive decoding process is accentuated. This problem is further exaggerated as CoTs scale in length in thinking models [GYZ+25] and as the number of repeated generations scale to elicit higher quality answers [BJE+24, WSL+24, SLXK24].
Figure 1: A full FF block will contain the complete performance of the original model without any benefit to efficiency. Sparse FF algorithms can be very efficient by using subsets of the FF block but harm math performance. Our method, Caprese, uses a sparse FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for highly performative inference in language and math settings while being efficient. Layers are drawn as trapezoids to highlight the relative size of inputs and outputs.
An ideal method should be efficient, performance preserving, and easy to integrate. Thankfully, low-rank structure in the feedforward (FF) output features can help make this possible. (We choose to focus on FF blocks since, contributing around 2/3 of an LLM’s parameters and over 50% of the generation latency.) Because of the success of works that exploit contextual sparsity in FF blocks on language tasks for efficiency, like CATS and GRIFFIN, we peek into the residuals of FF sparsity-based efficient methods. Using an oracle top-$k$ filter on FF nonlinearity output magnitudes, we observe huge reductions in error with a low-rank approximation to the FF output residuals (Figure 2). Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively mundane. This observation motivates us to estimate the residual from sparse FF methods with low-rank layers.
Figure 2: Average relative FF output error of generated tokens with varying top-$k$ densities and low-rank approximations. The density is the fraction of non-zero intermediate FF features maintained by top-$k$. Samples consist of 16 random MATH prompts and generations by the original model. A relatively small low-rank approximation to the top-$k$ residual dramatically reduces error much more effectively than just increasing density.
We introduce Caprese [^1] Our method’s components are modular with easy preparation like a Caprese salad. (CAPability REcovery with Scalable Efficiency) to learn FF residuals from any sparse FF algorithm using small low-rank linear layers for improved performance (Figure 1). Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome both challenges of math reasoning and makes significant progress towards an ideal efficient method:
  1. 1.
    Performance Enhancement: Demonstrated across a variety of models (e.g., instruct and thinking models), Caprese recovers much if not all of the math performance lost from deploying a sparse FF algorithm without harming language tasks. Also, Caprese maintains its performance benefit when applying different techniques of test-time scaling. For example, scaling generations with Caprese on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
  2. 2.
    Efficiency: Due to the ability to combine our low-rank layers with existing FF layers, Caprese reduces generation latency. For instance, Caprese reduces latency by 11.7% when generating 2048 tokens using DeepSeek-R1-Distill-Qwen 14B with GRIFFIN.
  3. 3.
    Low Budget: The distillation process of Caprese is cheap since we are able to see significant gains in math performance by training on only 20K synthetic math samples. Moreover, with a low-rank layer size of only 256, this equates to adding roughly $0.8\%$ additional parameters into Llama 3.1 8B and Gemma 2 9B, dwarfed by savings in active parameters ($\sim$2B cut for both models with Caprese (CATS)).
While Caprese is not restricted to math tasks, we focus on math in this paper, as current efficient sparse methods can obliterate an LLM’s math capabilities. Future work can investigate its benefit to other domains.
The remainder of this paper is organized as follows. Section 2 covers background on related works and describes GRIFFIN and CATS, two efficient sparse FF algorithms that serve as baselines. Section 3 details Caprese’s architecture and distillation procedure. Then, we showcase the strong performance of Caprese in Section 4: instruct LLM inference (Section 4.1), scaling generation outputs (Section 4.2), and scaling generation length with thinking LLMs (Section 4.3). Finally, we quantify latency improvements in Section 4.4.
2 Background
In this section, we provide an overview of the efficient LLM inference literature that Caprese builds upon with in-depth descriptions of GRIFFIN and CATS which we use as baselines and some recent developments for test-time scaling.
2.1 Related Works
Efficient Inference for FF Blocks.
To improve the efficiency of FF blocks in LLMs, various methods leverage existing sparse structures in FF features [GSBL20, DLBZ22, LYB+22, DCC23, LWD+23]. Pruning sets a parameter subset to zero to enforce static sparsity [LDS89, SLBK23, FA23, MFW23]. Mixtures of experts (MoEs) adaptively select predefined parameter subsets in FF blocks, though they tend to require significant fine-tuning or training from scratch [JJNH91, ZLL+21, DDZ+24, LWD+23]. Similar to MoEs, another line of work aims at exploiting contextual sparsity to dynamically select FF neurons for each input without training. GRIFFIN [DCC24] is a calibration-free method that uses FF activations from the prefill phase to adaptively prune FF neurons for generation. CATS [LLZ+24] uses hard thresholding to skip computation in parts of the FF block with a custom kernel. More details of GRIFFIN and CATS can be found in Section 2.2.
Shorter CoTs.
Several works aim to improve reasoning efficiency by directly reducing the number of generated tokens [RG24, NRS+24, AZ25, ABH25, XXZH25, QYS+25]. Since we aim to reduce the per-token latency, this line of research to encourage brevity is orthogonal to ours and could be used alongside our method.
Test-time Scaling.
To enhance LLM performance, the priority has historically been to scale training with more data and bigger models [KMH+20, HBM+22]. More recently, there is an increasing effort towards test-time scaling to boost performance on difficult tasks like math. Two ways of test-time scaling that have seen great success. First, for a single prompt, multiple responses can be sampled from the LLM [BJE+24, SLXK24, WSL+24, MSE24, SHZ+24]. This increases the probability that a desired response lies in the pool of responses. While this method of scaling is highly parallelizable, it also relies on a verifier model to select the best one. Second, CoTs can be lengthened for each response [GYZ+25]. By extrapolating the success of CoTs to extreme lengths (e.g., DeepSeek-R1 generates up to 32K tokens per response), the accuracy of the final answer improves significantly, but this method has low parallelizability due to the autoregressive nature of generation.
2.2 Training-free & Adaptively Sparse FF Methods
We now briefly describe the inner workings of GRIFFIN and CATS. Let $\bm{X}\in\mathbb{R}^{S\times D}$ be the input into the FF block during the prefill phase with sequence length $S$ and feature size $D$. Define the FF block as $\text{FF}(\bm{X})=\text{FF}_{2}(\text{FF}_{1}(\bm{X}))$ such that
$\bm{Z}=\text{FF}_{1}(\bm{X})$ $=\sigma(\bm{X}\bm{W}_{g})\odot\bm{X}\bm{W}_{1},$ (1)
$\text{FF}_{2}(\bm{Z})$ $=\bm{Z}\bm{W}_{2},$ (2)
where $\bm{W}_{g},\bm{W}_{1}\in\mathbb{R}^{D\times D_{\text{FF}}}$, $\bm{W}_{2}\in\mathbb{R}^{D_{\text{FF}}\times D}$, $\sigma$ is a nonlinear function, $\odot$ is an element-wise multiplication operator, and $D_{\text{FF}}$ is the FF intermediate feature size. Typically, $D_{\text{FF}}\gg D$. Although recent LLMs use this architecture, for LLMs without gated functions (e.g., OPT [ZRG+22]), $\bm{X}\bm{W}_{1}$ can be removed. Bias terms are omitted for brevity.
GRIFFIN.
GRIFFIN [DCC24] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns. GRIFFIN calculates $[\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $[\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$. The result $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ gives a metric to perform top-$k$ selection across corresponding columns of $\bm{W}_{g}$ and $\bm{W}_{1}$, and rows of $\bm{W}_{2}$ to produce $\widehat{\bm{W}}_{g},\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{k\times D}$. Then, for the generation phase, the following FF block is used for input $\bm{x}\in\mathbb{R}^{D}$:
$\widehat{\bm{z}}=\widehat{\text{FF}}_{1}(\bm{x})$ $=\sigma(\bm{x}\widehat{\bm{W}}_{g})\odot\bm{x}\widehat{\bm{W}}_{1},$ (3)
$\widehat{\text{FF}}_{2}(\widehat{\bm{z}})$ $=\widehat{\bm{z}}\widehat{\bm{W}}_{2}.$ (4)
The compressed FF blocks are fixed throughout generation but are dynamic across prompts.
CATS.
CATS [LLZ+24] uses hard thresholding to skip computation in part of the FF block. Letting $T_{\tau}$ be the hard thresholding function with threshold $\tau$, CATS computes
$\widehat{\bm{z}}=\widehat{\text{FF}}_{1}(\bm{x})$ $=T_{\tau}(\sigma(\bm{x}\bm{W}_{g}))\odot\bm{x}\bm{W}_{1},$ (5)
$\widehat{\text{FF}}_{2}(\widehat{\bm{z}})$ $=\widehat{\bm{z}}\bm{W}_{2}.$ (6)
The weights $\bm{W}_{1}$ and $\bm{W}_{2}$ should be sparsified into $\widehat{\bm{W}}_{1}$ and $\widehat{\bm{W}}_{2}$, respectively, based on the non-zero entries of $T_{\tau}(\sigma(\bm{x}\bm{W}_{g}))$ for latency improvement, which can vary from token to token and require a custom kernel for wall clock speed-up. The parameter $\tau$ is calibrated to be a desired percentile on a dataset. In this paper, to avoid calibration, we set $\tau$ based on prefill features and only threshold during the generation phase, analogous to GRIFFIN.
3 Method: Caprese
Motivated by the low-rank structure of residuals in Figure 2, we introduce Caprese which distills approximation errors in embeddings into low-rank linear layers in FF blocks. See Figure 1 for an illustration of our method.
3.1 Layer-wise Distillation
Inference efficiency algorithms often introduce feature approximation errors in favor of faster generation, which we mitigate with distillation. Let the efficient and approximate FF block be $\widehat{\text{FF}}(\bm{x})=\widehat{\text{FF}}_{2}(\widehat{\text{FF}}_{1}(\bm{x}))$. In our design of Caprese, we do not constrain $\widehat{\text{FF}}$ to be any specific method or architecture. For instance, $\widehat{\text{FF}}$ could be an FF block with GRIFFIN [DCC24] or CATS [LLZ+24]. Then, the error is
$\|\text{FF}(\bm{x})-\widehat{\text{FF}}(\bm{x})\|_{2}^{2}.$
We choose to reduce this residual with a low-rank linear layer, meaning we want to solve
$\min_{\bm{L},\bm{R}}\frac{1}{|\mathcal{X}|}\sum_{\bm{x}\in\mathcal{X}}\|\text{FF}(\bm{x})-\widehat{\text{FF}}(\bm{x})-\bm{x}\bm{L}\bm{R}\|_{2}^{2}$ (7)
for input set $\mathcal{X}$, $\bm{L}\in\mathbb{R}^{D\times r}$, and $\bm{R}\in\mathbb{R}^{r\times D}$ where $r\ll D_{\text{FF}}$. The optimal solution can be computed analytically since this is a reduced rank regression problem, but the size of $|\mathcal{X}|$ and $D$ may make it prohibitively expensive. Therefore, we opt to learn $\bm{L}$ and $\bm{R}$ independently for every FF block (i.e., previous FF blocks are assumed to be from the original model), allowing for parallel layer-wise training. This takes inspiration from LESS [DYZ+24] which uses layer-wise training on attention residuals for key-value cache compression. Each $\bm{R}$ is initialized as a zero matrix since the efficient approximation is assumed to be of good quality, and original model weights are frozen.
Even if the layer-wise MSE is minimized, this does not guarantee the final outputs closely resemble the original model due to error accumulation. Hence, we also distill end-to-end (E2E) to further improve the performance.
3.2 End-to-end Distillation
Using the learned low-rank layers as an initialization, we put them all together to distill the final model embedding before the linear head into the efficient model, again using MSE:
$\min_{(\bm{L}_{i},\bm{R}_{i})_{i=1,\dots,L}}\frac{1}{|\mathcal{X}|}\sum_{\bm{x}\in\mathcal{X}}\|\text{M}(\bm{x})-\text{M}_{\text{student}}(\bm{x})\|_{2}^{2}$ (8)
where M and $\text{M}_{\text{student}}$ are the original LLM and efficient LLM with distillation layers, respectively, excluding the final linear head. $L$ is the number of layers in the the model. All original weights are frozen, so the only tunable parameters are the $\bm{L}$ and $\bm{R}$ of each FF block.
3.3 Parallel Inference Computation
The computation of $\bm{x}\bm{L}\bm{R}$ can be done in parallel with the original FF operations. In fact, $\bm{L}$ and $\bm{R}$ can be concatenated with the up and down projection matrices, respectively. In other words, the Caprese FF block is $\widehat{\text{FF}}^{+}(\bm{x})=\widehat{\text{FF}}^{+}_{2}(\widehat{\text{FF}}^{+}_{1}(\bm{x}))$, such that
$\widehat{\bm{z}}^{+}=\widehat{\text{FF}}^{+}_{1}(\bm{x})$ $=\begin{bmatrix}\sigma(\bm{x}\widehat{\bm{W}}_{g})&\bm{1}_{r}\end{bmatrix}\odot\bm{x}\begin{bmatrix}\widehat{\bm{W}}_{1}&\bm{L}\end{bmatrix},$ (9)
$\widehat{\text{FF}}^{+}_{2}(\widehat{\bm{z}}^{+})$ $=\widehat{\bm{z}}^{+}\begin{bmatrix}\widehat{\bm{W}}_{2}^{\top}&\bm{R}^{\top}\end{bmatrix}^{\top},$ (10)
where $\bm{1}_{r}$ is a one-vector with length $r$. In practice, to save memory and time, we do not materialize $\bm{1}_{r}$ but directly assign the product with $\sigma(\bm{x}\widehat{\bm{W}}_{g})$ to corresponding entries of $\bm{x}\widehat{\bm{W}}^{+}_{1}$. Recall that the prefill stage still just uses the original model.
3.4 Training Details
We set the inner dimension of our low-rank layer to $r=256$ (to see the effect of different $r$, see Appendix A). In comparison to the enormous inner dimension of FF layers (e.g., $D_{\text{FF}}=14336$ for Llama 3.1 8B and Gemma 2 9B), our choice of $r$ is relatively miniscule, adding only $\sim 1\%$ new parameters for all tested models (Appendix B).
We use a 20K subset of a common synthetic math training set for training. For a fair comparison, the same subset is used for both layer-wise and E2E distillation for every model. Each training sample is prepended with a CoT instruction: “Please reason step by step.” At test time, the actual instructions may be vastly different. Layer-wise and E2E training consists of 20 epochs and 3 epochs, respectively. Training and inference are done in BF16. Hyperparameters are listed in Appendix E.
Model GSM8K MATH CoQA QASPER
Llama 3.1 8B Instruct 27.90 13.94 63.88 15.16
[…]
Layer-wise Caprese 27.14 9.72 65.05 14.21
E2E Caprese 40.49 12.64 65.50 15.35
[…]
Layer-wise Caprese 37.68 13.66 63.92 14.34
E2E Caprese 51.86 13.88 64.27 14.42
Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43
[…]
Layer-wise Caprese 13.72 6.62 56.07 13.40
E2E Caprese 21.00 8.44 56.55 13.88
[…]
Layer-wise Caprese 18.65 8.28 55.58 14.35
E2E Caprese 20.39 9.04 56.12 13.75
Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45
[…]
Layer-wise Caprese 40.18 13.70 64.33 11.60
E2E Caprese 44.66 16.96 64.83 12.35
[…]
Layer-wise Caprese 45.49 13.62 60.53 12.26
E2E Caprese 46.85 14.40 61.72 12.36
Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
[…]
Layer-wise Caprese 42.53 12.32 63.77 10.75
E2E Caprese 48.14 13.70 63.37 11.05
[…]
Layer-wise Caprese 46.32 13.90 61.92 10.82
E2E Caprese 46.17 14.16 63.92 11.03
Gemma 2 9B Instruct 78.17 27.64 63.78 9.91
[…]
Layer-wise Caprese 76.72 25.84 64.20 9.82
E2E Caprese 76.65 25.30 64.42 9.92
[…]
Layer-wise Caprese 77.18 28.16 64.52 10.46
E2E Caprese 77.18 28.00 64.87 10.02
Table 1: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER). FF sparsity is set to 50%, and $r=256$.
4 Experiments
We showcase the effectiveness of Caprese at recovering much, if not all, of the math performance lost from efficient inference algorithms without sacrificing efficiency or performance on language tasks. In the following section (Section 4.1), we observe better math performance for instruct LLMs without losing their generalizability. Then, we study Caprese’s benefit on two axes of test-time scaling: number of generations (Section 4.2) and generation length with thinking models (Section 4.3). Finally, we highlight Caprese’s latency improvements in Section 4.4. When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying sparse method and similarly for GRIFFIN. Otherwise, we use “Caprese” for brevity. Unless specified, the FF intermediate feature sparsity is set at 50%. NOTE: For a more meaningful baseline, we only apply GRIFFIN to the first half of the model for all experiments since we observed much steeper drops in math performance if used for all layers. Regardless, Caprese still improves the performance of GRIFFIN on all layers. CATS is applied to all layers.
4.1 Instruct Models
We first looking into different instruct LLMs which are performative on a variety of non-reasoning tasks and decent reasoning performance. We show Caprese is able to preserve math performance without sacrificing performance on pure language tasks like question answering. In more detail, we test on the Llama 3 family [DJP+24] and Gemma 2 family [TRP+24] of models on zero-shot GSM8K [CKB+21], MATH [HBK+21], CoQA [RCM19], and QASPER [DLB+21], using LM Evaluation Harness [GTA+23] and CoT prompts for math tasks.
Table 1 shows that Caprese is able to preserve most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math. In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese is able to effectively recover the lost performance. With the exception of Gemma 2 9B Instruct where the layer-wise distillation appears to be better, E2E Caprese is the most performative in the majority of math scenarios, suggesting a benefit beyond minimizing local error. Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A). The best performers in CoQA and QASPER are a toss-up, but all methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases). CATS and Caprese considerably improve GSM8K performance of Llama 3.1 8B Instruct, which has trouble following the prompt template with zero-shot.
4.2 Scaling Best-of-$N$: Coverage
Figure 3: Coverage and standard deviation of 140 samples from MATH as the number of generation attempts, $N$, scales. We define Relative Compute $\text{Relative Compute Units}=N\times A$ where $A$ is the fraction of total parameters activated per input. FF sparsity set to 50%. Best viewed zoomed.
Now, we see the generalizability of Caprese on the first axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$ (BoN). Increasing the number of generations per prompt, $N$, is one way to scale inference compute for improved performance, as the probability of generating a correct answer increases. To evaluate, we find the coverage (Pass@$K$) on 140 samples from MATH. We use an oracle verifier to accurately assess the quality of the pool of generated responses. To combat the high variance when $K$ is close to $N$, we calculate the average coverage for $K=1,\dots,100$ across 10 independent pools of 100 generations for each sample. Factoring in the saved compute, E2E Caprese is able to have similar or better coverage scaling compared to the original model, as shown in Figure 3. Notably, E2E Caprese (GRIFFIN) on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
4.3 Scaling Length: Thinking Models
Model MATH-500 AIME 2024 AMC 2023
DeepSeek-R1-Distill-Qwen 1.5B 79.40 31.11 $\pm$ 1.92 57.50 $\pm$ 5.00
[…]
Layer-wise Caprese 47.20 3.33 $\pm$ 0.00 27.50 $\pm$ 2.50
E2E Caprese 60.40 6.67 $\pm$ 3.33 35.83 $\pm$ 7.22
[…]
Layer-wise Caprese 73.80 16.67 $\pm$ 3.33 46.67 $\pm$ 2.89
E2E Caprese 74.80 21.11 $\pm$ 1.92 48.33 $\pm$ 3.82
DeepSeek-R1-Distill-Qwen 7B 90.20 51.11 $\pm$ 3.85 76.67 $\pm$ 2.89
[…]
Layer-wise Caprese 84.80 27.78 $\pm$ 1.92 60.83 $\pm$ 1.44
E2E Caprese 85.40 28.89 $\pm$ 6.94 70.00 $\pm$ 5.00
[…]
Layer-wise Caprese 87.00 35.56 $\pm$ 5.09 65.83 $\pm$ 7.22
E2E Caprese 90.00 32.22 $\pm$ 5.09 79.17 $\pm$ 5.77
DeepSeek-R1-Distill-Qwen 14B 92.80 63.33 $\pm$ 3.33 90.83 $\pm$ 2.89
[…]
Layer-wise Caprese 90.80 38.89 $\pm$ 1.92 87.50 $\pm$ 5.00
E2E Caprese 89.20 37.78 $\pm$ 10.72 82.50 $\pm$ 6.61
[…]
Layer-wise Caprese 91.00 60.00 $\pm$ 5.77 88.33 $\pm$ 1.44
E2E Caprese 92.00 58.89 $\pm$ 11.71 85.83 $\pm$ 3.82
Table 2: Thinking models’ 0-shot accuracies on math reasoning tasks. Sparsity is set at 50%, and $r=256$. AIME 2024 and AMC 2023 columns also include the sample standard deviation across 3 runs. Further improvements with reselection are shown in Table 3.
Thinking models provide another axis of test-time scaling by augmenting the CoT length before giving a final answer. Because this entails very long generation lengths, it is critical that error accumulation across tokens be minimized. We test on DeepSeek-R1-Distill-Qwen models [YYZ+24, GYZ+25] using the prompt templates and configurations from Open R1 [Fac25] (temperature and top-$p$ set to 0.6 and 0.95, respectively). For these models, we evaluate zero-shot performance on MATH-500 [HBK+21, LKB+23], AIME 2024, and AMC 2023 for a max generation length of 32768. Due to the small number of problems in AIME 2024 (30 problems) and AMC 2023 (40 problems), we evaluate these 3 independent times, reporting the average accuracy and sample standard deviation.
From Table 2, Caprese is the most performative in most cases. All methods generally appear to be more robust as the model size increases. For instance, CATS and GRIFFIN had very little impact on DeepSeek-R1-Distill-Qwen 14B’s accuracy on MATH-500, with Caprese performing similarly. By far, AIME 2024 is the most challenging and noisy task (due to its difficult and small dataset), with severe degradation from CATS and GRIFFIN and partial recovery using our method. In contrast, AMC 2023 results are less noisy and regained performance from our method is much more substantial. In the next section, we show a simple way to push this recovery even further with reselection.
4.3.1 Enhanced Performance with Reselection
We can push the performance of Caprese with neuron reselection. For a sample, GRIFFIN and CATS calculate metrics ($\bm{s}$ and $\tau$) to determine subsets of the FF block to use, but these metrics are fixed during the generation phase. After generating many tokens, these selection metrics can benefit from regular updates mid-generation.
For GRIFFIN, updating the metric $\bm{s}$ entails integrating the FF feature statistics of generated tokens into $\bm{s}$. While this can be done by re-running prefill on all tokens, there is a more efficient way. This can be done by passing in the generated tokens following the last reselection through the full model. As these tokens propagate through each layer, we find the selection metric for the generated tokens $\bm{s}_{G}$ and update corresponding KV pairs. This is similar to verification in speculative decoding [CBI+23, LKM23]. Since $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we can define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use. This setup updates the pruned layers yet avoids prefill for all tokens. Table 3 shows a clear benefit of reselection (even if infrequent) in Caprese by pushing the performance much closer to the full model.
Model No Reselect $\rho=1024$ $\rho=256$ Full
DeepSeek-R1-Distill-Qwen 1.5B 35.83 $\pm$ 7.22 40.83 $\pm$ 5.77 40.00 $\pm$ 7.50 57.50 $\pm$ 5.00
DeepSeek-R1-Distill-Qwen 7B 70.00 $\pm$ 5.00 68.17 $\pm$ 1.44 74.17 $\pm$ 5.20 76.67 $\pm$ 3.82
DeepSeek-R1-Distill-Qwen 14B 82.50 $\pm$ 6.61 87.50 $\pm$ 4.33 89.17 $\pm$ 3.82 90.83 $\pm$ 2.89
Table 3: E2E Caprese AMC 2023 accuracies and standard deviations when recalculating the GRIFFIN pruning metric during generation. $\rho$ is the rate at which we reselect pruned neurons in terms of generated tokens. No reselection and the full model are special cases where $\rho=\infty$ and $\rho=1$, respectively.
Reselection is also possible with CATS but requires recomputing prefill for all tokens. The parameter to update is the hard thresholding parameter $\tau$, but because this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory. Consequently, prefill will need to be redone anytime we want to update $\tau$. Additionally, since $\tau$ represents a cutoff for a desired percentile (e.g., median), it is fairly robust to new observations. Given this and the lack of computational incentive, we focus primarily on reselection for GRIFFIN.
4.3.2 Effect on Natural Response Length
Figure 4: Average number of response tokens for MATH-500 queries across different problem difficulties (top) and subjects (bottom). The subjects are algebra, counting & probability, geometry, intermediate algebra, number theory, prealgebra, and precalculus. The global averages are indicated by the dashed lines. Sparsity is set at 50%.
Caprese implicitly encourages brevity, often even producing shorter responses than from the full thinking model. Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference (except for cutting off generation at 32K tokens). Shown in Figure 4 with MATH-500, the shortest mean response length for all problem difficulties and subjects is either the full model or Caprese. Meanwhile, CATS consistently outputs the longest responses, averaging roughly 1K more across every sample. It is also interesting to note that with increasing problem difficulty, all models and methods naturally allocate more inference tokens towards answering the question. See Appendix C for similar observations with GRIFFIN.
Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DeepSeek-R1-Distill-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit. As we will see in the next section, we can also reduce the per-token latency for further computational savings.
4.4 Efficiency
Caprese reduces end-to-end generation latency. Table 4 shows the latencies of different setups and models. Caprese cuts latency by >11% when generating 2048 or fewer tokens from a prompt of 2048 tokens using DeepSeek-R1-Distill-Qwen 14B (Qwen 2.5 14B architecture) or Gemma 2 9B. Latency improvements are smaller with generation length 8192 but stilln faster than the full model. Moreover, the latency differences between GRIFFIN and Caprese are fairly small, suggesting that Caprese has minimal overhead. Since we use GRIFFIN on half of each model, the latency reductions are smaller than [DCC24] recorded. Metrics were collected on an NVIDIA L40 GPU using BF16 precision and pre-allocated memory for the KV cache. Time to first/next token metrics are included in Appendix D.
Model Setup Full GRIFFIN Caprese
Qwen 2.5 14B 2048+256 12.8 11.1 ($-13.3\%$) 11.3 ($-11.7\%$)
2048+2048 113.3 98.6 ($-13.0\%$) 100.1 ($-11.7\%$)
2048+8192 661.9 603.2 ($-8.9\%$) 603.7 ($-8.8\%$)
Gemma 2 9B 2048+256 9.6 8.4 ($-12.5\%$) 8.5 ($-11.5\%$)
2048+2048 84.4 74.5 ($-11.7\%$) 75.1 ($-11.0\%$)
2048+8192 483.6 444.4 ($-8.1\%$) 445.8 ($-7.8\%$)
Table 4: End-to-end generation latency (s) for GRIFFIN and Caprese (GRIFFIN). For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively. Percentages in parentheses show the reduction in latency. As before, GRIFFIN is applied to the first half of the model, sparsity is set at 50%, and $r=256$.
5 Conclusion
To combat the inefficiency of the long and brittle generation process associated with math reasoning, we introduce Caprese, a highly performant and efficient method with strong math and language capabilities that is compatible with a broad class of efficient FF algorithms. In the future, it would be interesting to evaluate its benefit in other reasoning domains besides math, explore the use of input-dependent low-rank layers, and investigate the knowledge contained in these low-rank layers. Caprese, along with future developments, pushes towards the long-term goal of degradation-free efficient LLM inference.
Acknowledgements
The work of H. Dong is supported in part by the Wei Shen and Xuehong Zhang Presidential Fellowship at Carnegie Mellon University.
Appendix A Effect of Rank & Sparsity
Here, we ablate the relationship between varying ranks in Caprese and sparsity levels in CATS. Using Llama 3.2 1B Instruct, we show the test performance of MATH in Figure 5. The same training procedure and data are used as outlined in Section 3. For all ablated ranks, Caprese consistently outperforms pure CATS by a large margin when CATS performs poorly relative to the full model. With a couple of exceptions (perhaps due to randomness in the generation process), greater performance is correlated with higher rank.
Figure 5: Llama 3.2 1B Instruct’s performance on MATH with varying densities of CATS and ranks in Caprese with end-to-end training.
Appendix B Caprese Parameters
Caprese has a tiny parameter footprint. In Figure 6, we see that Caprese adds roughly 1% new parameters relative to the full model with a trend downwards as model size increases.
Figure 6: The percent of new parameters that Caprese ($r=256$) adds relative to the entire model (left) and relative to only the FF parameters (right) for the Llama 3, Gemma 2, and Qwen 2.5 model families.
Appendix C GRIFFIN Response Lengths
Complementing Figure 4 for CATS, we show the natural response lengths to MATH-500 samples in Figure 7. Here, we see the similar observations as in Section 4.3.2, though the difference between Caprese (GRIFFIN) and the full model is slightly larger but decreasing with model size. Even so, response lengths of Caprese is still significantly shorter than GRIFFIN.
Figure 7: Average number of response tokens for MATH-500 queries across different problem difficulties (top) and subjects (bottom) when using GRIFFIN. The subjects are algebra, counting & probability, geometry, intermediate algebra, number theory, prealgebra, and precalculus. The global averages are indicated by the dashed lines. Sparsity is set at 50%.
Appendix D Additional Efficiency Metrics
Using the same setup from Section 4.4, we also report the average time to first token (TTFT) and time to next token (TTNT) in Table 5. The GRIFFIN and Caprese add a small overhead during the prefill phase but generally improve TTNT by 7ms and 5ms for Qwen 2.5 14B and Gemma 2 9B, respectively. There is hardly any difference in TTFT and TTNT between GRIFFIN and Caprese, reinforcing the fact that Caprese adds negligible overhead on top of GRIFFIN.
Model Setup Avg. Time to First Token Avg. Time to Next Token
Full GRIFFIN Caprese Full GRIFFIN Caprese
Qwen 2.5 14B 2048+256 0.64 0.67 0.67 0.048 0.041 0.042
2048+2048 0.67 0.68 0.68 0.055 0.048 0.049
2048+8192 0.77 0.80 0.80 0.081 0.074 0.074
Gemma 2 9B 2048+256 0.42 0.46 0.46 0.036 0.031 0.032
2048+2048 0.45 0.49 0.50 0.041 0.036 0.036
2048+8192 0.50 0.53 0.53 0.059 0.054 0.054
Table 5: Time to first and next token (s) for GRIFFIN and Caprese (GRIFFIN). For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively. As before, GRIFFIN is applied to the first half of the model, sparsity is set at 50%, and $r=256$.
Appendix E Training Hyperparameters
Table 6 lists the hyperparameter settings for training Caprese layers. The E2E learning rates lie in the interval [4e-6, 2e-4], where larger models tend to learn better with smaller learning rates.
Layer-wise End-to-end
Optimizer Adam Adam
Learning rate 1e-3 [4e-6, 2e-4] (varies)
Batch size 128 16
Epochs 20 3
Training samples 2e5 2e5
Scheduler Linear Linear
Warmup 2% 2%
Table 6: Caprese layer-wise and E2E training hyperparameters.

### v2

@@
 With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the math capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs.
-Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>11% reduction to generate 2048 tokens with Qwen 2.5 14B) while encouraging response brevity.
+Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>16% time-to-next-token reduction) while encouraging response brevity (up to 8.5% fewer tokens).
 1 Introduction
@@
 These tasks tend to demand long generation which drives up latency, making efficiency a dire issue.
-Fortunately, many efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization.
-However, for math reasoning tasks, many efficient inference algorithms begin to break down, decimating performance, despite their robustness in language settings.
+Fortunately, many sparsity-based efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization.
+However, for math reasoning tasks, many of these algorithms begin to break down, decimating performance, despite their robustness in language settings.
 Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and math capabilities.
 One of the main differences between math reasoning and many language tasks is the generation length.
-Reasoning typically involves long generation due to improved performance from chain-of-thought (CoT) [WWS+22], which often surpasses the length of the input query.
+Reasoning typically involves long generations from chain-of-thoughts (CoTs) [WWS+22], which often far surpass the length of the input query.
 However, CoT performance falls apart with efficient algorithms that introduce approximation errors which build up over time.
@@
 Similarly, for thinking models: applying GRIFFIN [DCC24], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [GYZ+25] drops MATH-500 [LKB+23] accuracy from 79.40% to 42.00% (Table 2).
-Moreover, since generation is much more demanding computationally than prefill (e.g., >150$\times$ slower to generate than prefill 1K tokens on one GPU with Llama 3.1 8B), there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling.
-Thus, this poses two key inference challenges with math reasoning:
-  1. 1.
-    Instability: Long CoTs can be sensitive to mistakes.
-Even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [ZCX+24].
-Hence, efficient inference algorithms should respect the delicacy of CoT.
-  2. 2.
-    Inefficiency: Due to the reliance on CoT, which tend to lead to long generation, the already computationally expensive LLM autoregressive decoding process is accentuated.
+Moreover, since generation is much more computationally demanding than prefill, there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling.
+Thus, this poses two key inference challenges with math reasoning.
+First, even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [ZCX+24].
+Second, the already computationally expensive LLM autoregressive decoding process is accentuated with generating CoTs.
 This problem is further exaggerated as CoTs scale in length in thinking models [GYZ+25] and as the number of repeated generations scale to elicit higher quality answers [BJE+24, WSL+24, SLXK24].
-Figure 1: A full FF block will contain the complete performance of the original model without any benefit to efficiency.
+An ideal method should be efficient, performance preserving, and easy to integrate.
+Figure 1: A full FF block maximizes accuracy without any benefit to efficiency.
 Sparse FF algorithms can be very efficient by using subsets of the FF block but harm math performance.
-Our method, Caprese, uses a sparse FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for highly performative inference in language and math settings while being efficient.
+Our method, Caprese, uses a sparse FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for performative inference in language and math settings while being efficient.
 Layers are drawn as trapezoids to highlight the relative size of inputs and outputs.
-An ideal method should be efficient, performance preserving, and easy to integrate.
 Thankfully, low-rank structure in the feedforward (FF) output features can help make this possible.
-(We choose to focus on FF blocks since, contributing around 2/3 of an LLM’s parameters and over 50% of the generation latency.) Because of the success of works that exploit contextual sparsity in FF blocks on language tasks for efficiency, like CATS and GRIFFIN, we peek into the residuals of FF sparsity-based efficient methods.
+(We choose to focus on FF blocks since they contribute around 2/3 of an LLM’s parameters and about 50% of the generation latency.) Because of the success of works that exploit contextual sparsity in FF blocks on language tasks for efficiency, like CATS and GRIFFIN, we peek into the residuals of FF sparsity-based efficient methods.
 Using an oracle top-$k$ filter on FF nonlinearity output magnitudes, we observe huge reductions in error with a low-rank approximation to the FF output residuals (Figure 2).
-Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively mundane.
+Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively tiny.
 This observation motivates us to estimate the residual from sparse FF methods with low-rank layers.
@@
 The density is the fraction of non-zero intermediate FF features maintained by top-$k$.
-Samples consist of 16 random MATH prompts and generations by the original model.
-A relatively small low-rank approximation to the top-$k$ residual dramatically reduces error much more effectively than just increasing density.
-We introduce Caprese [^1] Our method’s components are modular with easy preparation like a Caprese salad.
-(CAPability REcovery with Scalable Efficiency) to learn FF residuals from any sparse FF algorithm using small low-rank linear layers for improved performance (Figure 1).
-Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome both challenges of math reasoning and makes significant progress towards an ideal efficient method:
+Samples consist of 16 random MATH generations by the original model.
+A relatively small low-rank approximation to the top-$k$ residual reduces error more effectively than increasing density.
+We introduce Caprese (CAPability REcovery with Scalable Efficiency) to learn FF residuals from any sparse FF algorithm using small low-rank linear layers for improved performance (Figure 1).
+While Caprese works for any task, we focus on math in this paper, as current efficient sparse methods can obliterate an LLM’s math capabilities.
+Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome the aforementioned challenges of math reasoning and makes significant progress towards an ideal efficient method:
   1. 1.
-    Performance Enhancement: Demonstrated across a variety of models (e.g., instruct and thinking models), Caprese recovers much if not all of the math performance lost from deploying a sparse FF algorithm without harming language tasks.
+    Performance Enhancement: Demonstrated across a variety of instruct and thinking models, Caprese recovers much if not all of the math performance lost from deploying a sparse FF algorithm without harming language tasks.
 Also, Caprese maintains its performance benefit when applying different techniques of test-time scaling.
-For example, scaling generations with Caprese on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
   2. 2.
-    Efficiency: Due to the ability to combine our low-rank layers with existing FF layers, Caprese reduces generation latency.
-For instance, Caprese reduces latency by 11.7% when generating 2048 tokens using DeepSeek-R1-Distill-Qwen 14B with GRIFFIN.
+    Efficiency: Due to its parallelizability with existing FF layers, Caprese adds negligible overhead, preserving latency reductions of the underlying method.
   3. 3.
@@
 Moreover, with a low-rank layer size of only 256, this equates to adding roughly $0.8\%$ additional parameters into Llama 3.1 8B and Gemma 2 9B, dwarfed by savings in active parameters ($\sim$2B cut for both models with Caprese (CATS)).
-While Caprese is not restricted to math tasks, we focus on math in this paper, as current efficient sparse methods can obliterate an LLM’s math capabilities.
-Future work can investigate its benefit to other domains.
-The remainder of this paper is organized as follows.
+Our extensive experiments on Caprese indicate strong performance on various tasks and models.
+For example, DeepSeek-R1-Distill-Qwen 7B with a sparse method drops AMC 2023 accuracy from 75.00% to 62.50%, but Caprese is able to bring it to 78.13%, beyond the original accuracy.
+Furthermore, scaling generation quantity with Caprese on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% fewer active parameters.
+We also show Caprese reduces latency by 16% using the Qwen 2.5 14B architecture.
+Paper Organization.
 Section 2 covers background on related works and describes GRIFFIN and CATS, two efficient sparse FF algorithms that serve as baselines.
@@
 Then, we showcase the strong performance of Caprese in Section 4: instruct LLM inference (Section 4.1), scaling generation outputs (Section 4.2), and scaling generation length with thinking LLMs (Section 4.3).
-Finally, we quantify latency improvements in Section 4.4.
+Finally, we quantify efficiency improvements in Section 4.4.
 2 Background
@@
 CATS [LLZ+24] uses hard thresholding to skip computation in parts of the FF block with a custom kernel.
-More details of GRIFFIN and CATS can be found in Section 2.2.
 Shorter CoTs.
@@
 While this method of scaling is highly parallelizable, it also relies on a verifier model to select the best one.
-Second, CoTs can be lengthened for each response [GYZ+25].
+Second, CoTs can be lengthened for each response [GYZ+25, MYS+25, YLY+25].
 By extrapolating the success of CoTs to extreme lengths (e.g., DeepSeek-R1 generates up to 32K tokens per response), the accuracy of the final answer improves significantly, but this method has low parallelizability due to the autoregressive nature of generation.
@@
 GRIFFIN [DCC24] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns.
-GRIFFIN calculates $[\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $[\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$.
+GRIFFIN calculates $\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$.
 The result $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ gives a metric to perform top-$k$ selection across corresponding columns of $\bm{W}_{g}$ and $\bm{W}_{1}$, and rows of $\bm{W}_{2}$ to produce $\widehat{\bm{W}}_{g},\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{k\times D}$.
@@
 $\widehat{\bm{z}}=\widehat{\text{FF}}_{1}(\bm{x})$ $=\sigma(\bm{x}\widehat{\bm{W}}_{g})\odot\bm{x}\widehat{\bm{W}}_{1},$ (3)
-$\widehat{\text{FF}}_{2}(\widehat{\bm{z}})$ $=\widehat{\bm{z}}\widehat{\bm{W}}_{2}.$ (4)
+[…]
 The compressed FF blocks are fixed throughout generation but are dynamic across prompts.
@@
 $\widehat{\bm{z}}=\widehat{\text{FF}}_{1}(\bm{x})$ $=T_{\tau}(\sigma(\bm{x}\bm{W}_{g}))\odot\bm{x}\bm{W}_{1},$ (5)
-$\widehat{\text{FF}}_{2}(\widehat{\bm{z}})$ $=\widehat{\bm{z}}\bm{W}_{2}.$ (6)
+[…]
 The weights $\bm{W}_{1}$ and $\bm{W}_{2}$ should be sparsified into $\widehat{\bm{W}}_{1}$ and $\widehat{\bm{W}}_{2}$, respectively, based on the non-zero entries of $T_{\tau}(\sigma(\bm{x}\bm{W}_{g}))$ for latency improvement, which can vary from token to token and require a custom kernel for wall clock speed-up.
@@
 Each $\bm{R}$ is initialized as a zero matrix since the efficient approximation is assumed to be of good quality, and original model weights are frozen.
-Even if the layer-wise MSE is minimized, this does not guarantee the final outputs closely resemble the original model due to error accumulation.
-Hence, we also distill end-to-end (E2E) to further improve the performance.
+We also distill end-to-end (E2E) to further improve the performance.
 3.2 End-to-end Distillation
@@
 where M and $\text{M}_{\text{student}}$ are the original LLM and efficient LLM with distillation layers, respectively, excluding the final linear head.
-$L$ is the number of layers in the the model.
+$L$ is the number of layers in the model.
 All original weights are frozen, so the only tunable parameters are the $\bm{L}$ and $\bm{R}$ of each FF block.
@@
 We set the inner dimension of our low-rank layer to $r=256$ (to see the effect of different $r$, see Appendix A).
-In comparison to the enormous inner dimension of FF layers (e.g., $D_{\text{FF}}=14336$ for Llama 3.1 8B and Gemma 2 9B), our choice of $r$ is relatively miniscule, adding only $\sim 1\%$ new parameters for all tested models (Appendix B).
+In comparison to the large inner dimension of FF layers (e.g., $D_{\text{FF}}=14336$ for Llama 3.1 8B and Gemma 2 9B), our choice of $r$ is relatively miniscule, adding only $\sim 1\%$ new parameters for all tested models (Appendix B).
 We use a 20K subset of a common synthetic math training set for training.
@@
 Hyperparameters are listed in Appendix E.
+3.5 Comparison with LoRA
+Our method derives inspiration from and shares a slight connection with low-rank adaptation (LoRA), a widely used method for efficient fine tuning with the addition of low-rank parameters [HSW+22], but remains distinct since they target different inefficiencies.
+To illustrate, recall a sparse algorithm constructs $\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$ by selecting columns from $\bm{W}_{1}$.
+Alternatively, construct $\widetilde{\bm{W}}_{1}\in\mathbb{R}^{D\times D_{\text{FF}}}$ by setting unwanted columns to zero, so nonzero columns in $\widetilde{\bm{W}}_{1}$ match columns in $\widehat{\bm{W}}_{1}$ and vice versa.
+Doing the same to columns of $\bm{W}_{g}$ and rows of $\bm{W}_{2}$, LoRA learns $\bm{A}_{1},\bm{A}_{g},\bm{B}_{2}^{\top}\in\mathbb{R}^{D\times r}$ and $\bm{B}_{1},\bm{B}_{g},\bm{A}_{2}^{\top}\in\mathbb{R}^{r\times D_{\text{FF}}}$ to form a new FF block with LoRA weights:
+$\widetilde{\text{FF}}(\bm{x})=\left(\sigma\big(\bm{x}(\widetilde{\bm{W}}_{g}+\bm{A}_{g}\bm{B}_{g})\big)\odot\big(\bm{x}(\widetilde{\bm{W}}_{1}+\bm{A}_{1}\bm{B}_{1})\big)\right)(\widetilde{\bm{W}}_{2}+\bm{A}_{2}\bm{B}_{2}).$ (11)
+From this, we see LoRA is not designed to be applied to sparse efficient inference algorithms.
+First, this is because in its compressed form, LoRA requires two sequential operations per linear layer.
+In contrast, Caprese parameters can be appended to existing ones, so additional computation is all parallel to the original operations (Section 3.3).
+The real latency impact of this difference is quantified later in Table 4.
+Caprese’s architecture also allows efficient training, but that is merely a byproduct of our design, not its purpose.
+Second, for a method like GRIFFIN which reduces the intermediate FF feature size from $D_{\text{FF}}$ to $k$, adding LoRA will negate this benefit while Caprese will preserve the reduced feature size.
+Due to this difference, we de-prioritize comparisons with LoRA in the main text, but include comparisons with LoRA in Appendix C for interested readers.
 Model GSM8K MATH CoQA QASPER
 Llama 3.1 8B Instruct 27.90 13.94 63.88 15.16
-[…]
+GRIFFIN 17.44 6.16 63.37 12.63
 Layer-wise Caprese 27.14 9.72 65.05 14.21
 E2E Caprese 40.49 12.64 65.50 15.35
-[…]
+CATS 40.56 11.80 58.85 12.26
 Layer-wise Caprese 37.68 13.66 63.92 14.34
@@
 Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43
-[…]
+GRIFFIN 7.13 5.42 56.05 14.11
 Layer-wise Caprese 13.72 6.62 56.07 13.40
 E2E Caprese 21.00 8.44 56.55 13.88
-[…]
+CATS 19.18 7.54 54.40 13.88
 Layer-wise Caprese 18.65 8.28 55.58 14.35
@@
 Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45
-[…]
+GRIFFIN 28.96 10.98 64.52 12.52
 Layer-wise Caprese 40.18 13.70 64.33 11.60
 E2E Caprese 44.66 16.96 64.83 12.35
-[…]
+CATS 41.24 12.04 58.87 11.36
 Layer-wise Caprese 45.49 13.62 60.53 12.26
@@
 Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
-[…]
+GRIFFIN 33.74 11.32 63.28 11.07
 Layer-wise Caprese 42.53 12.32 63.77 10.75
 E2E Caprese 48.14 13.70 63.37 11.05
-[…]
+CATS 34.42 10.56 61.53 10.11
 Layer-wise Caprese 46.32 13.90 61.92 10.82
@@
 Gemma 2 9B Instruct 78.17 27.64 63.78 9.91
-[…]
+GRIFFIN 59.21 25.22 63.82 10.14
 Layer-wise Caprese 76.72 25.84 64.20 9.82
 E2E Caprese 76.65 25.30 64.42 9.92
-[…]
+CATS 76.50 27.32 64.37 9.78
 Layer-wise Caprese 77.18 28.16 64.52 10.46
 E2E Caprese 77.18 28.00 64.87 10.02
-Table 1: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER). FF sparsity is set to 50%, and $r=256$.
+Table 1: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER). $F$ sparsity is set to 50%, and $r=256$.
 4 Experiments
@@
 Then, we study Caprese’s benefit on two axes of test-time scaling: number of generations (Section 4.2) and generation length with thinking models (Section 4.3).
-Finally, we highlight Caprese’s latency improvements in Section 4.4.
+Finally, we highlight Caprese’s latency and length improvements in Section 4.4.
 When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying sparse method and similarly for GRIFFIN.
@@
 NOTE: For a more meaningful baseline, we only apply GRIFFIN to the first half of the model for all experiments since we observed much steeper drops in math performance if used for all layers.
-Regardless, Caprese still improves the performance of GRIFFIN on all layers.
 CATS is applied to all layers.
 4.1 Instruct Models
-We first looking into different instruct LLMs which are performative on a variety of non-reasoning tasks and decent reasoning performance.
-We show Caprese is able to preserve math performance without sacrificing performance on pure language tasks like question answering.
-In more detail, we test on the Llama 3 family [DJP+24] and Gemma 2 family [TRP+24] of models on zero-shot GSM8K [CKB+21], MATH [HBK+21], CoQA [RCM19], and QASPER [DLB+21], using LM Evaluation Harness [GTA+23] and CoT prompts for math tasks.
+We first look into different instruct LLMs which are performative on non-reasoning tasks and have decent reasoning performance.
+We show Caprese is able to preserve math performance without sacrificing quality on pure language tasks like question answering.
+We test on the Llama 3 family [DJP+24] and Gemma 2 family [TRP+24] of models on 0-shot GSM8K [CKB+21], MATH [HBK+21], CoQA [RCM19], and QASPER [DLB+21], using CoT prompts for math tasks.
 Table 1 shows that Caprese is able to preserve most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math.
 In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese is able to effectively recover the lost performance.
-With the exception of Gemma 2 9B Instruct where the layer-wise distillation appears to be better, E2E Caprese is the most performative in the majority of math scenarios, suggesting a benefit beyond minimizing local error.
+E2E Caprese is the most performative in the majority of math scenarios, suggesting a benefit beyond minimizing local error.
 Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A).
-The best performers in CoQA and QASPER are a toss-up, but all methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases).
+The best performers for CoQA and QASPER are a toss-up, but all methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases).
 CATS and Caprese considerably improve GSM8K performance of Llama 3.1 8B Instruct, which has trouble following the prompt template with zero-shot.
 4.2 Scaling Best-of-$N$: Coverage
-Figure 3: Coverage and standard deviation of 140 samples from MATH as the number of generation attempts, $N$, scales.
-We define Relative Compute $\text{Relative Compute Units}=N\times A$ where $A$ is the fraction of total parameters activated per input.
-FF sparsity set to 50%.
+Figure 3: Coverage and standard deviation of 140 samples from MATH as the number of generation attempts, NN, scales.
+We define $\text{Relative Compute Units}=N\times A$ where $A$ is the fraction of total parameters activated per input.
+$F$ sparsity set to 50%.
 Best viewed zoomed.
-Now, we see the generalizability of Caprese on the first axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$ (BoN).
+Now, we see the generalizability of Caprese on the first axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$.
 Increasing the number of generations per prompt, $N$, is one way to scale inference compute for improved performance, as the probability of generating a correct answer increases.
@@
 4.3 Scaling Length: Thinking Models
-Model MATH-500 AIME 2024 AMC 2023
-DeepSeek-R1-Distill-Qwen 1.5B 79.40 31.11 $\pm$ 1.92 57.50 $\pm$ 5.00
-[…]
-Layer-wise Caprese 47.20 3.33 $\pm$ 0.00 27.50 $\pm$ 2.50
-E2E Caprese 60.40 6.67 $\pm$ 3.33 35.83 $\pm$ 7.22
-[…]
-Layer-wise Caprese 73.80 16.67 $\pm$ 3.33 46.67 $\pm$ 2.89
-E2E Caprese 74.80 21.11 $\pm$ 1.92 48.33 $\pm$ 3.82
-DeepSeek-R1-Distill-Qwen 7B 90.20 51.11 $\pm$ 3.85 76.67 $\pm$ 2.89
-[…]
-Layer-wise Caprese 84.80 27.78 $\pm$ 1.92 60.83 $\pm$ 1.44
-E2E Caprese 85.40 28.89 $\pm$ 6.94 70.00 $\pm$ 5.00
-[…]
-Layer-wise Caprese 87.00 35.56 $\pm$ 5.09 65.83 $\pm$ 7.22
-E2E Caprese 90.00 32.22 $\pm$ 5.09 79.17 $\pm$ 5.77
-DeepSeek-R1-Distill-Qwen 14B 92.80 63.33 $\pm$ 3.33 90.83 $\pm$ 2.89
-[…]
-Layer-wise Caprese 90.80 38.89 $\pm$ 1.92 87.50 $\pm$ 5.00
-E2E Caprese 89.20 37.78 $\pm$ 10.72 82.50 $\pm$ 6.61
-[…]
-Layer-wise Caprese 91.00 60.00 $\pm$ 5.77 88.33 $\pm$ 1.44
-E2E Caprese 92.00 58.89 $\pm$ 11.71 85.83 $\pm$ 3.82
+Model MATH-500 AIME 2024 AMC 2023 BRUMO 2025
+DS-Qwen 1.5B 79.40 30.00 $\pm$ 2.72 60.83 $\pm$ 7.47 25.83 $\pm$ 3.19
+GRIFFIN 42.00 2.50 $\pm$ 1.67 16.88 $\pm$ 1.25 1.67 $\pm$ 1.92
+Layer-wise Caprese 47.20 2.50 $\pm$ 1.67 28.13 $\pm$ 2.39 0.00 $\pm$ 0.00
+E2E Caprese 60.40 6.67 $\pm$ 2.72 34.38 $\pm$ 6.57 4.17 $\pm$ 1.67
+CATS 72.00 11.67 $\pm$ 6.94 37.50 $\pm$ 3.54 10.83 $\pm$ 4.19
+Layer-wise Caprese 73.80 15.83 $\pm$ 3.19 48.13 $\pm$ 3.75 13.33 $\pm$ 2.72
+E2E Caprese 74.80 20.83 $\pm$ 1.67 47.50 $\pm$ 3.54 20.00 $\pm$ 2.72
+DS-Qwen 7B 90.20 47.50 $\pm$ 7.87 75.00 $\pm$ 4.08 45.83 $\pm$ 4.19
+GRIFFIN 80.60 21.67 $\pm$ 4.30 62.50 $\pm$ 8.90 25.00 $\pm$ 8.39
+Layer-wise Caprese 84.80 29.17 $\pm$ 3.19 64.38 $\pm$ 7.18 27.50 $\pm$ 4.19
+E2E Caprese 85.40 30.00 $\pm$ 6.09 71.88 $\pm$ 5.54 29.17 $\pm$ 5.69
+CATS 89.20 34.17 $\pm$ 7.88 62.50 $\pm$ 2.04 37.50 $\pm$ 7.39
+Layer-wise Caprese 87.00 35.00 $\pm$ 4.30 70.00 $\pm$ 10.21 39.17 $\pm$ 3.19
+E2E Caprese 90.00 33.33 $\pm$ 4.71 78.13 $\pm$ 5.15 47.50 $\pm$ 5.00
+DS-Qwen 14B 92.80 63.33 $\pm$ 2.72 90.63 $\pm$ 2.39 60.83 $\pm$ 4.19
+GRIFFIN 89.80 32.50 $\pm$ 7.39 80.63 $\pm$ 3.75 33.33 $\pm$ 3.85
+Layer-wise Caprese 90.80 41.67 $\pm$ 5.77 86.88 $\pm$ 4.27 43.33 $\pm$ 4.71
+E2E Caprese 89.20 39.17 $\pm$ 9.18 84.38 $\pm$ 6.57 40.00 $\pm$ 2.72
+CATS 92.80 58.34 $\pm$ 1.92 88.75 $\pm$ 1.44 55.00 $\pm$ 3.33
+Layer-wise Caprese 91.00 61.67 $\pm$ 5.77 88.75 $\pm$ 1.44 53.00 $\pm$ 7.20
+E2E Caprese 92.00 58.34 $\pm$ 9.62 87.50 $\pm$ 2.04 49.17 $\pm$ 4.19
 Table 2: Thinking models’ 0-shot accuracies on math reasoning tasks.
 Sparsity is set at 50%, and $r=256$.
-AIME 2024 and AMC 2023 columns also include the sample standard deviation across 3 runs.
+AIME 2024 and AMC 2023 columns also include the sample standard deviation across 4 runs.
 Further improvements with reselection are shown in Table 3.
@@
 Because this entails very long generation lengths, it is critical that error accumulation across tokens be minimized.
-We test on DeepSeek-R1-Distill-Qwen models [YYZ+24, GYZ+25] using the prompt templates and configurations from Open R1 [Fac25] (temperature and top-$p$ set to 0.6 and 0.95, respectively).
-For these models, we evaluate zero-shot performance on MATH-500 [HBK+21, LKB+23], AIME 2024, and AMC 2023 for a max generation length of 32768.
-Due to the small number of problems in AIME 2024 (30 problems) and AMC 2023 (40 problems), we evaluate these 3 independent times, reporting the average accuracy and sample standard deviation.
+We test on DeepSeek-R1-Distill-Qwen (which we abbreviate to DS-Qwen) models [YYZ+24, GYZ+25] using the prompt templates and configurations from Open R1 [Fac25] (temperature and top-$p$ set to 0.6 and 0.95, respectively).
+For these models, we test 0-shot performance on MATH-500 [HBK+21, LKB+23], AIME 2024 [AIM25], AMC 2023 [AMC23], and BRUMO 2025 [BRU25] for a max generation length of 32768.
+Due to the small size of AIME 2024, AMC 2023, and BRUMO 2025, we evaluate them 4 times, reporting the average accuracy and sample standard deviation.
 From Table 2, Caprese is the most performative in most cases.
-All methods generally appear to be more robust as the model size increases.
-For instance, CATS and GRIFFIN had very little impact on DeepSeek-R1-Distill-Qwen 14B’s accuracy on MATH-500, with Caprese performing similarly.
-By far, AIME 2024 is the most challenging and noisy task (due to its difficult and small dataset), with severe degradation from CATS and GRIFFIN and partial recovery using our method.
-In contrast, AMC 2023 results are less noisy and regained performance from our method is much more substantial.
+DS-Qwen 1.5B and 7B see the greatest benefit with E2E Caprese usually achieving the highest accuracy, sometimes even exceeding the full model’s performance.
+DS-Qwen 7B with CATS brings AMC 2023 accuracy down to 62.50% from 75.00%, but Caprese lifts performance above the original model’s to 78.13%, and similarly with DS-Qwen 7B with CATS and BRUMO 2025.
+All methods are more robust as the model size increases.
+For instance, CATS and GRIFFIN had very little impact on DS-Qwen 14B’s accuracy on MATH-500, with Caprese performing similarly.
+AIME 2024 and BRUMO 2025 challenge CATS and GRIFFIN with several degradation and partial recovery using our method.
 In the next section, we show a simple way to push this recovery even further with reselection.
@@
 We can push the performance of Caprese with neuron reselection.
-For a sample, GRIFFIN and CATS calculate metrics ($\bm{s}$ and $\tau$) to determine subsets of the FF block to use, but these metrics are fixed during the generation phase.
-After generating many tokens, these selection metrics can benefit from regular updates mid-generation.
+For a sample, GRIFFIN and CATS calculate metrics ($\bm{s}$ and $\tau$) to determine subsets of the FF block to use, but these metrics are fixed during generation.
+Updating them mid-generation can benefit downstream performance.
 For GRIFFIN, updating the metric $\bm{s}$ entails integrating the FF feature statistics of generated tokens into $\bm{s}$.
-While this can be done by re-running prefill on all tokens, there is a more efficient way.
-This can be done by passing in the generated tokens following the last reselection through the full model.
+While this can be done by re-running prefill on all tokens, there is a more efficient way by passing in the generated tokens following the last reselection through the full model.
 As these tokens propagate through each layer, we find the selection metric for the generated tokens $\bm{s}_{G}$ and update corresponding KV pairs.
-This is similar to verification in speculative decoding [CBI+23, LKM23].
-Since $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we can define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use.
-This setup updates the pruned layers yet avoids prefill for all tokens.
-Table 3 shows a clear benefit of reselection (even if infrequent) in Caprese by pushing the performance much closer to the full model.
-Model No Reselect $\rho=1024$ $\rho=256$ Full
-DeepSeek-R1-Distill-Qwen 1.5B 35.83 $\pm$ 7.22 40.83 $\pm$ 5.77 40.00 $\pm$ 7.50 57.50 $\pm$ 5.00
-DeepSeek-R1-Distill-Qwen 7B 70.00 $\pm$ 5.00 68.17 $\pm$ 1.44 74.17 $\pm$ 5.20 76.67 $\pm$ 3.82
-DeepSeek-R1-Distill-Qwen 14B 82.50 $\pm$ 6.61 87.50 $\pm$ 4.33 89.17 $\pm$ 3.82 90.83 $\pm$ 2.89
-Table 3: E2E Caprese AMC 2023 accuracies and standard deviations when recalculating the GRIFFIN pruning metric during generation.
-$\rho$ is the rate at which we reselect pruned neurons in terms of generated tokens.
-No reselection and the full model are special cases where $\rho=\infty$ and $\rho=1$, respectively.
+This is like verification in speculative decoding [CBI+23, LKM23].
+As $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use (as described in Section 2).
+This updates the pruned layers yet avoids prefill for all tokens.
+Table 3 shows a clear benefit of reselection (even if infrequent compared to speculative decoding) in Caprese by pushing the performance much closer to the full model’s as we decrease the number of decode steps between reselection rounds.
+Although GRIFFIN is generally more harmful to accuracy than CATS due to its structured pruning, paired with Caprese and reselection, Caprese (GRIFFIN) can exceed the performance of Caprese (CATS) and reach the full model performance.
+Model No Reselect $\rho=1024$ $\rho=256$ $\rho=64$ Full
+DS-Qwen 1.5B 34.38 41.87 45.00 58.75 60.83
+DS-Qwen 7B 71.88 69.25 73.13 73.75 75.00
+DS-Qwen 14B 84.38 88.13 88.75 91.88 90.63
+Table 3: E2E Caprese AMC 2023 accuracies and standard deviations when recalculating GRIFFIN pruning metrics and reselecting pruned neurons every $\rho$ decode steps.
+No reselection and the full model are special cases where $\rho=\infty$ and $\rho=0$, respectively.
 Reselection is also possible with CATS but requires recomputing prefill for all tokens.
 The parameter to update is the hard thresholding parameter $\tau$, but because this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory.
-Consequently, prefill will need to be redone anytime we want to update $\tau$.
-Additionally, since $\tau$ represents a cutoff for a desired percentile (e.g., median), it is fairly robust to new observations.
+In turn, prefill will need to be redone anytime we want to update $\tau$.
+Since $\tau$ represents a cutoff for a desired percentile (e.g., median), it is fairly robust to new observations.
 Given this and the lack of computational incentive, we focus primarily on reselection for GRIFFIN.
-4.3.2 Effect on Natural Response Length
-Figure 4: Average number of response tokens for MATH-500 queries across different problem difficulties (top) and subjects (bottom).
-The subjects are algebra, counting & probability, geometry, intermediate algebra, number theory, prealgebra, and precalculus.
-The global averages are indicated by the dashed lines.
-Sparsity is set at 50%.
+4.4 Efficiency
+Caprese reduces generation latency.
+Table 4 shows the latencies of different generation settings.
+Caprese cuts total latency and time to next token by >16% for the Qwen 2.5 14B architecture.
+Moreover, the latency differences between GRIFFIN and Caprese are exceptionally small, suggesting that Caprese has minimal overhead.
+We also include latencies of GRIFFIN paired with LoRA weights, which although faster than the base model, negates >40% of the time savings that GRIFFIN has provided.
+This highlights the efficiency suboptimality of LoRA during inference.
+Metrics were collected on an NVIDIA L40 GPU using BF16 precision.
+Setup End-to-end Latency (s) Avg. Time to Next Token (ms)
+Full GRIFFIN LoRA Caprese Full GRIFFIN LoRA Caprese
+2048+256 10.5 8.7 9.5 8.7 41 34 37 34
+2048+2048 84.4 70.3 76.5 70.4 41 34 37 34
+2048+8192 344.9 287.7 312.5 288.8 42 35 38 35
+Table 4: End-to-end generation latency (s) and average time to next token (ms) for GRIFFIN, LoRA, and Caprese using Qwen 2.5 14B.
+For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively.
+As before, GRIFFIN is applied to the first half of the model, sparsity is 50%, and $r=256$.
+4.4.1 Bonus: Effect on Natural Response Length
+Figure 4: Average number of response tokens for MATH-500 queries with increasing problem difficulty. The global averages are indicated by the dashed lines. Sparsity is set at 50%.
 Caprese implicitly encourages brevity, often even producing shorter responses than from the full thinking model.
 Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference (except for cutting off generation at 32K tokens).
-Shown in Figure 4 with MATH-500, the shortest mean response length for all problem difficulties and subjects is either the full model or Caprese.
+Shown in Figure 4 with MATH-500, the shortest mean response length for all problem difficulties is either the full model or Caprese.
 Meanwhile, CATS consistently outputs the longest responses, averaging roughly 1K more across every sample.
 It is also interesting to note that with increasing problem difficulty, all models and methods naturally allocate more inference tokens towards answering the question.
-See Appendix C for similar observations with GRIFFIN.
-Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DeepSeek-R1-Distill-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit.
-As we will see in the next section, we can also reduce the per-token latency for further computational savings.
-4.4 Efficiency
-Caprese reduces end-to-end generation latency.
-Table 4 shows the latencies of different setups and models.
-Caprese cuts latency by >11% when generating 2048 or fewer tokens from a prompt of 2048 tokens using DeepSeek-R1-Distill-Qwen 14B (Qwen 2.5 14B architecture) or Gemma 2 9B.
-Latency improvements are smaller with generation length 8192 but stilln faster than the full model.
-Moreover, the latency differences between GRIFFIN and Caprese are fairly small, suggesting that Caprese has minimal overhead.
-Since we use GRIFFIN on half of each model, the latency reductions are smaller than [DCC24] recorded.
-Metrics were collected on an NVIDIA L40 GPU using BF16 precision and pre-allocated memory for the KV cache.
-Time to first/next token metrics are included in Appendix D.
-Model Setup Full GRIFFIN Caprese
-Qwen 2.5 14B 2048+256 12.8 11.1 ($-13.3\%$) 11.3 ($-11.7\%$)
-2048+2048 113.3 98.6 ($-13.0\%$) 100.1 ($-11.7\%$)
-2048+8192 661.9 603.2 ($-8.9\%$) 603.7 ($-8.8\%$)
-Gemma 2 9B 2048+256 9.6 8.4 ($-12.5\%$) 8.5 ($-11.5\%$)
-2048+2048 84.4 74.5 ($-11.7\%$) 75.1 ($-11.0\%$)
-2048+8192 483.6 444.4 ($-8.1\%$) 445.8 ($-7.8\%$)
-Table 4: End-to-end generation latency (s) for GRIFFIN and Caprese (GRIFFIN).
-For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively.
-Percentages in parentheses show the reduction in latency.
-As before, GRIFFIN is applied to the first half of the model, sparsity is set at 50%, and $r=256$.
+Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DS-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit.
+This brevity of Caprese is more pronounced in larger models, proportionally, from a 5.8% reduction in tokens compared to the full DS-Qwen 7B model to an 8.5% reduction compared to the full DS-Qwen 14B model.
+The cause of this phenomenon is of interest for future work.
+See Appendix D for similar observations with GRIFFIN.
 5 Conclusion
@@
 Figure 6: The percent of new parameters that Caprese ($r=256$) adds relative to the entire model (left) and relative to only the FF parameters (right) for the Llama 3, Gemma 2, and Qwen 2.5 model families.
-Appendix C GRIFFIN Response Lengths
+Appendix C LoRA Performance
+We report the performance of using LoRA in place of Caprese while keeping parameter count constant in Table 5.
+Caprese achieves the highest accuracy in slightly more cases than LoRA, but as per our discussion on LoRA in Section 3.5 and efficiency analysis in Section 4.4, LoRA’s main benefit is efficient training, not efficient inference.
+Model GSM8K MATH CoQA QASPER
+Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43
+GRIFFIN 7.13 5.42 56.05 14.11
+Layer-wise LoRA 8.11 5.60 56.00 14.65
+E2E LoRA 20.47 10.10 56.10 14.41
+Layer-wise Caprese 13.72 6.62 56.07 13.40
+E2E Caprese 21.00 8.44 56.55 13.88
+Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45
+GRIFFIN 28.96 10.98 64.52 12.52
+Layer-wise LoRA 26.46 11.28 63.88 12.55
+E2E LoRA 42.91 16.50 64.67 12.63
+Layer-wise Caprese 40.18 13.70 64.33 11.60
+E2E Caprese 44.66 16.96 64.83 12.35
+Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
+GRIFFIN 33.74 11.32 63.28 11.07
+Layer-wise LoRA 39.58 14.66 62.98 10.54
+E2E LoRA 44.66 16.71 63.03 10.81
+Layer-wise Caprese 42.53 12.32 63.77 10.75
+E2E Caprese 48.14 13.70 63.37 11.05
+Table 5: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER) with LoRA.
+$F$ sparsity is set to 50%, $r=256$, and LoRA ranks are set to match the number of parameters that Caprese adds.
+Appendix D GRIFFIN Response Lengths
 Complementing Figure 4 for CATS, we show the natural response lengths to MATH-500 samples in Figure 7.
@@
 Even so, response lengths of Caprese is still significantly shorter than GRIFFIN.
-Figure 7: Average number of response tokens for MATH-500 queries across different problem difficulties (top) and subjects (bottom) when using GRIFFIN.
-The subjects are algebra, counting & probability, geometry, intermediate algebra, number theory, prealgebra, and precalculus.
-The global averages are indicated by the dashed lines.
-Sparsity is set at 50%.
-Appendix D Additional Efficiency Metrics
-Using the same setup from Section 4.4, we also report the average time to first token (TTFT) and time to next token (TTNT) in Table 5.
-The GRIFFIN and Caprese add a small overhead during the prefill phase but generally improve TTNT by 7ms and 5ms for Qwen 2.5 14B and Gemma 2 9B, respectively.
-There is hardly any difference in TTFT and TTNT between GRIFFIN and Caprese, reinforcing the fact that Caprese adds negligible overhead on top of GRIFFIN.
-Model Setup Avg. Time to First Token Avg. Time to Next Token
-Full GRIFFIN Caprese Full GRIFFIN Caprese
-Qwen 2.5 14B 2048+256 0.64 0.67 0.67 0.048 0.041 0.042
-2048+2048 0.67 0.68 0.68 0.055 0.048 0.049
-2048+8192 0.77 0.80 0.80 0.081 0.074 0.074
-Gemma 2 9B 2048+256 0.42 0.46 0.46 0.036 0.031 0.032
-2048+2048 0.45 0.49 0.50 0.041 0.036 0.036
-2048+8192 0.50 0.53 0.53 0.059 0.054 0.054
-Table 5: Time to first and next token (s) for GRIFFIN and Caprese (GRIFFIN).
-For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively.
-As before, GRIFFIN is applied to the first half of the model, sparsity is set at 50%, and $r=256$.
+Figure 7: Average number of response tokens for MATH-500 queries with increasing problem difficulty. The global averages are indicated by the dashed lines. Sparsity is set at 50%.
 Appendix E Training Hyperparameters

### v3

@@
-Scalable LLM Math Reasoning Acceleration with Low-rank Distillation
+Scalable LLM Reasoning Acceleration with Low-rank Distillation
 Abstract
@@
 In this paper, we propose Caprese, a resource-efficient distillation method to recover lost capabilities from deploying efficient inference methods, focused primarily in feedforward blocks.
-With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the math capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs.
+With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the reasoning capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs.
 Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>16% time-to-next-token reduction) while encouraging response brevity (up to 8.5% fewer tokens).
 1 Introduction
-With the increasing capabilities of large language models (LLMs) [VSP+17], evaluations are also becoming increasingly sophisticated, typically involving multi-step reasoning such as in math problem solving.
+With the increasing capabilities of large language models (LLMs) [52], outputs are also becoming increasingly sophisticated, typically involving multi-step reasoning such as in math problem solving.
 These tasks tend to demand long generation which drives up latency, making efficiency a dire issue.
 Fortunately, many sparsity-based efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization.
-However, for math reasoning tasks, many of these algorithms begin to break down, decimating performance, despite their robustness in language settings.
-Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and math capabilities.
-One of the main differences between math reasoning and many language tasks is the generation length.
-Reasoning typically involves long generations from chain-of-thoughts (CoTs) [WWS+22], which often far surpass the length of the input query.
+However, for reasoning tasks, many of these algorithms begin to break down, decimating performance, despite their robustness in language settings.
+Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and reasoning capabilities.
+One of the main differences between reasoning and many language tasks is the generation length.
+Reasoning typically involves long generations from chain-of-thoughts (CoTs) [53], which often far surpass the length of the input query.
 However, CoT performance falls apart with efficient algorithms that introduce approximation errors which build up over time.
-For instance, deploying CATS [LLZ+24], a sparse thresholding method, on Gemma 2 2B [TRP+24] has little impact on language generation performance, but knocks GSM8K [CKB+21] accuracy from 51.02% to 34.42% (Table 1).
-Similarly, for thinking models: applying GRIFFIN [DCC24], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [GYZ+25] drops MATH-500 [LKB+23] accuracy from 79.40% to 42.00% (Table 2).
+For instance, deploying CATS [31], a sparse thresholding method, on Gemma 2 2B [51] has little impact on language generation performance, but knocks GSM8K [8] accuracy from 51.02% to 34.42% (Table 4).
+Similarly, for thinking models: applying GRIFFIN [16], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [23] drops MATH-500 [34] accuracy from 79.40% to 42.00% (Table 1).
 Moreover, since generation is much more computationally demanding than prefill, there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling.
 Thus, this poses two key inference challenges with math reasoning.
-First, even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [ZCX+24].
+First, even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [60].
 Second, the already computationally expensive LLM autoregressive decoding process is accentuated with generating CoTs.
-This problem is further exaggerated as CoTs scale in length in thinking models [GYZ+25] and as the number of repeated generations scale to elicit higher quality answers [BJE+24, WSL+24, SLXK24].
+This problem is further exaggerated as CoTs scale in length in thinking models [23] and as the number of repeated generations scale to elicit higher quality answers [5, 54, 48].
 An ideal method should be efficient, performance preserving, and easy to integrate.
 Figure 1: A full FF block maximizes accuracy without any benefit to efficiency.
-Sparse FF algorithms can be very efficient by using subsets of the FF block but harm math performance.
-Our method, Caprese, uses a sparse FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for performative inference in language and math settings while being efficient.
-Layers are drawn as trapezoids to highlight the relative size of inputs and outputs.
+Compressed $F$ algorithms can be very efficient by using subsets of the FF block but harm math performance.
+Our method, Caprese, uses a compressed FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for performative inference in language and math settings while being efficient.
+Layers are drawn as trapezoids to highlight the expansion of the intermediate feature size compared to the input size.
 Thankfully, low-rank structure in the feedforward (FF) output features can help make this possible.
@@
 Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively tiny.
-This observation motivates us to estimate the residual from sparse FF methods with low-rank layers.
+This observation motivates us to estimate the residual from FF compression methods with low-rank layers.
 Figure 2: Average relative FF output error of generated tokens with varying top-$k$ densities and low-rank approximations.
@@
 A relatively small low-rank approximation to the top-$k$ residual reduces error more effectively than increasing density.
-We introduce Caprese (CAPability REcovery with Scalable Efficiency) to learn FF residuals from any sparse FF algorithm using small low-rank linear layers for improved performance (Figure 1).
-While Caprese works for any task, we focus on math in this paper, as current efficient sparse methods can obliterate an LLM’s math capabilities.
-Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome the aforementioned challenges of math reasoning and makes significant progress towards an ideal efficient method:
+We introduce Caprese (CAPability REcovery with Scalable Efficiency) to learn FF residuals from a compressed FF algorithm using small low-rank linear layers for improved performance (Figure 1).
+While Caprese works for any task, we focus on reasoning (e.g., math) in this paper, as current efficient compression methods can obliterate an LLM’s reasoning capabilities.
+Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome the aforementioned challenges of reasoning and makes significant progress towards an ideal efficient method:
   1. 1.
-    Performance Enhancement: Demonstrated across a variety of instruct and thinking models, Caprese recovers much if not all of the math performance lost from deploying a sparse FF algorithm without harming language tasks.
+    Performance Enhancement: Demonstrated across a variety of instruct and thinking models, Caprese recovers much if not all of the reasoning performance lost from deploying a compressed FF algorithm without harming language tasks.
 Also, Caprese maintains its performance benefit when applying different techniques of test-time scaling.
@@
   3. 3.
-    Low Budget: The distillation process of Caprese is cheap since we are able to see significant gains in math performance by training on only 20K synthetic math samples.
+    Low Budget: The distillation process of Caprese is cheap since we are able to see significant gains in performance by training on only 20K synthetic math samples.
 Moreover, with a low-rank layer size of only 256, this equates to adding roughly $0.8\%$ additional parameters into Llama 3.1 8B and Gemma 2 9B, dwarfed by savings in active parameters ($\sim$2B cut for both models with Caprese (CATS)).
@@
 Paper Organization.
-Section 2 covers background on related works and describes GRIFFIN and CATS, two efficient sparse FF algorithms that serve as baselines.
+Section 2 covers related works and describes GRIFFIN and CATS, two efficient sparse FF compression algorithms that serve as baselines.
 Section 3 details Caprese’s architecture and distillation procedure.
-Then, we showcase the strong performance of Caprese in Section 4: instruct LLM inference (Section 4.1), scaling generation outputs (Section 4.2), and scaling generation length with thinking LLMs (Section 4.3).
+Then, we showcase the strong performance of Caprese in Section 4: scaling generation length with thinking LLMs (Section 4.1), scaling generation outputs (Section 4.2), and instruct LLM inference (Section 4.3).
 Finally, we quantify efficiency improvements in Section 4.4.
 2 Background
-In this section, we provide an overview of the efficient LLM inference literature that Caprese builds upon with in-depth descriptions of GRIFFIN and CATS which we use as baselines and some recent developments for test-time scaling.
+Here, we provide an overview of the efficient LLM inference literature that Caprese builds upon with in-depth descriptions of GRIFFIN and CATS which we use as baselines.
 2.1 Related Works
 Efficient Inference for FF Blocks.
-To improve the efficiency of FF blocks in LLMs, various methods leverage existing sparse structures in FF features [GSBL20, DLBZ22, LYB+22, DCC23, LWD+23].
-Pruning sets a parameter subset to zero to enforce static sparsity [LDS89, SLBK23, FA23, MFW23].
-Mixtures of experts (MoEs) adaptively select predefined parameter subsets in FF blocks, though they tend to require significant fine-tuning or training from scratch [JJNH91, ZLL+21, DDZ+24, LWD+23].
+To improve the efficiency of FF blocks in LLMs, various methods leverage existing sparse structures in FF features [22, 12, 33, 15, 36, 46].
+Pruning sets a parameter subset to zero to enforce static sparsity [30, 50, 21, 37].
+Mixtures of experts (MoEs) adaptively select predefined parameter subsets in FF blocks, though they tend to require significant fine-tuning or training from scratch [28, 59, 9, 36].
 Similar to MoEs, another line of work aims at exploiting contextual sparsity to dynamically select FF neurons for each input without training.
-GRIFFIN [DCC24] is a calibration-free method that uses FF activations from the prefill phase to adaptively prune FF neurons for generation.
-CATS [LLZ+24] uses hard thresholding to skip computation in parts of the FF block with a custom kernel.
+GRIFFIN [16] is a calibration-free method that uses FF activations from the prefill phase to adaptively prune FF neurons for generation.
+CATS [31] uses hard thresholding to skip computation in parts of the FF block with a custom kernel.
+Though not constrained to FF blocks, quantization is another popular method for compression [13, 11, 17, 35].
 Shorter CoTs.
-Several works aim to improve reasoning efficiency by directly reducing the number of generated tokens [RG24, NRS+24, AZ25, ABH25, XXZH25, QYS+25].
+Several works aim to improve reasoning efficiency by directly reducing the number of generated tokens [45, 41, 3, 4, 55, 42].
 Since we aim to reduce the per-token latency, this line of research to encourage brevity is orthogonal to ours and could be used alongside our method.
 Test-time Scaling.
-To enhance LLM performance, the priority has historically been to scale training with more data and bigger models [KMH+20, HBM+22].
+To enhance LLM performance, the priority has historically been to scale training with more data and bigger models [29, 26].
 More recently, there is an increasing effort towards test-time scaling to boost performance on difficult tasks like math.
 Two ways of test-time scaling that have seen great success.
-First, for a single prompt, multiple responses can be sampled from the LLM [BJE+24, SLXK24, WSL+24, MSE24, SHZ+24].
+First, for a single prompt, multiple responses can be sampled from the LLM [5, 48, 54, 38, 49, 14].
 This increases the probability that a desired response lies in the pool of responses.
 While this method of scaling is highly parallelizable, it also relies on a verifier model to select the best one.
-Second, CoTs can be lengthened for each response [GYZ+25, MYS+25, YLY+25].
+Second, CoTs can be lengthened for each response [23, 39, 56].
 By extrapolating the success of CoTs to extreme lengths (e.g., DeepSeek-R1 generates up to 32K tokens per response), the accuracy of the final answer improves significantly, but this method has low parallelizability due to the autoregressive nature of generation.
-2.2 Training-free & Adaptively Sparse FF Methods
+2.2 Training-free & Adaptively Sparse FF Compression Methods
 We now briefly describe the inner workings of GRIFFIN and CATS.
@@
 Typically, $D_{\text{FF}}\gg D$.
-Although recent LLMs use this architecture, for LLMs without gated functions (e.g., OPT [ZRG+22]), $\bm{X}\bm{W}_{1}$ can be removed.
+Although recent LLMs use this architecture, for LLMs without gated functions (e.g., OPT [58]), $\bm{X}\bm{W}_{1}$ can be removed.
 Bias terms are omitted for brevity.
 GRIFFIN.
-GRIFFIN [DCC24] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns.
+GRIFFIN [16] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns.
 GRIFFIN calculates $\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$.
@@
 CATS.
-CATS [LLZ+24] uses hard thresholding to skip computation in part of the FF block. Letting $T_{\tau}$ be the hard thresholding function with threshold $\tau$, CATS computes
+CATS [31] uses hard thresholding to skip computation in part of the FF block. Letting $T_{\tau}$ be the hard thresholding function with threshold $\tau$, CATS computes
 $\widehat{\bm{z}}=\widehat{\text{FF}}_{1}(\bm{x})$ $=T_{\tau}(\sigma(\bm{x}\bm{W}_{g}))\odot\bm{x}\bm{W}_{1},$ (5)
@@
 In our design of Caprese, we do not constrain $\widehat{\text{FF}}$ to be any specific method or architecture.
-For instance, $\widehat{\text{FF}}$ could be an FF block with GRIFFIN [DCC24] or CATS [LLZ+24].
+For instance, $\widehat{\text{FF}}$ could be an FF block with GRIFFIN [16], CATS [31], or quantized weights.
 Then, the error is
@@
 Therefore, we opt to learn $\bm{L}$ and $\bm{R}$ independently for every FF block (i.e., previous FF blocks are assumed to be from the original model), allowing for parallel layer-wise training.
-This takes inspiration from LESS [DYZ+24] which uses layer-wise training on attention residuals for key-value cache compression.
+This takes inspiration from LESS [18] which uses layer-wise training on attention residuals for key-value cache compression.
 Each $\bm{R}$ is initialized as a zero matrix since the efficient approximation is assumed to be of good quality, and original model weights are frozen.
@@
 Training and inference are done in BF16.
-Hyperparameters are listed in Appendix E.
+Hyperparameters are listed in Appendix G.
 3.5 Comparison with LoRA
-Our method derives inspiration from and shares a slight connection with low-rank adaptation (LoRA), a widely used method for efficient fine tuning with the addition of low-rank parameters [HSW+22], but remains distinct since they target different inefficiencies.
+Our method derives inspiration from and shares a slight connection with low-rank adaptation (LoRA), a widely used method for efficient fine tuning with the addition of low-rank parameters [27], but remains distinct since they target different inefficiencies.
 To illustrate, recall a sparse algorithm constructs $\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$ by selecting columns from $\bm{W}_{1}$.
@@
 Due to this difference, we de-prioritize comparisons with LoRA in the main text, but include comparisons with LoRA in Appendix C for interested readers.
-Model GSM8K MATH CoQA QASPER
-Llama 3.1 8B Instruct 27.90 13.94 63.88 15.16
-GRIFFIN 17.44 6.16 63.37 12.63
-Layer-wise Caprese 27.14 9.72 65.05 14.21
-E2E Caprese 40.49 12.64 65.50 15.35
-CATS 40.56 11.80 58.85 12.26
-Layer-wise Caprese 37.68 13.66 63.92 14.34
-E2E Caprese 51.86 13.88 64.27 14.42
-Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43
-GRIFFIN 7.13 5.42 56.05 14.11
-Layer-wise Caprese 13.72 6.62 56.07 13.40
-E2E Caprese 21.00 8.44 56.55 13.88
-CATS 19.18 7.54 54.40 13.88
-Layer-wise Caprese 18.65 8.28 55.58 14.35
-E2E Caprese 20.39 9.04 56.12 13.75
-Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45
-GRIFFIN 28.96 10.98 64.52 12.52
-Layer-wise Caprese 40.18 13.70 64.33 11.60
-E2E Caprese 44.66 16.96 64.83 12.35
-CATS 41.24 12.04 58.87 11.36
-Layer-wise Caprese 45.49 13.62 60.53 12.26
-E2E Caprese 46.85 14.40 61.72 12.36
-Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
-GRIFFIN 33.74 11.32 63.28 11.07
-Layer-wise Caprese 42.53 12.32 63.77 10.75
-E2E Caprese 48.14 13.70 63.37 11.05
-CATS 34.42 10.56 61.53 10.11
-Layer-wise Caprese 46.32 13.90 61.92 10.82
-E2E Caprese 46.17 14.16 63.92 11.03
-Gemma 2 9B Instruct 78.17 27.64 63.78 9.91
-GRIFFIN 59.21 25.22 63.82 10.14
-Layer-wise Caprese 76.72 25.84 64.20 9.82
-E2E Caprese 76.65 25.30 64.42 9.92
-CATS 76.50 27.32 64.37 9.78
-Layer-wise Caprese 77.18 28.16 64.52 10.46
-E2E Caprese 77.18 28.00 64.87 10.02
-Table 1: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER). $F$ sparsity is set to 50%, and $r=256$.
+Table 1: Thinking models’ 0-shot accuracies on reasoning tasks.
+Sparsity is set at 50%, and $r=256$.
+AIME 2024, AMC 2023, and BRUMO 2025 columns also include the sample standard deviation across 4 runs.
+Further improvements with reselection are shown in Table 3.
+Model MATH-500 AIME 2024 AMC 2023 BRUMO 2025 GPQA
+DS-Qwen 1.5B 79.40 30.00 $\pm$ 2.72 60.83 $\pm$ 7.47 25.83 $\pm$ 3.19 18.69
+GRIFFIN 42.00 2.50 $\pm$ 1.67 16.88 $\pm$ 1.25 1.67 $\pm$ 1.92 11.62
+Layer-wise Caprese 47.20 2.50 $\pm$ 1.67 28.13 $\pm$ 2.39 0.00 $\pm$ 0.00 13.13
+E2E Caprese 60.40 6.67 $\pm$ 2.72 34.38 $\pm$ 6.57 4.17 $\pm$ 1.67 16.67
+CATS 72.00 11.67 $\pm$ 6.94 37.50 $\pm$ 3.54 10.83 $\pm$ 4.19 11.62
+Layer-wise Caprese 73.80 15.83 $\pm$ 3.19 48.13 $\pm$ 3.75 13.33 $\pm$ 2.72 16.67
+E2E Caprese 74.80 20.83 $\pm$ 1.67 47.50 $\pm$ 3.54 20.00 $\pm$ 2.72 22.22
+DS-Qwen 7B 90.20 47.50 $\pm$ 7.87 75.00 $\pm$ 4.08 45.83 $\pm$ 4.19 38.38
+GRIFFIN 80.60 21.67 $\pm$ 4.30 62.50 $\pm$ 8.90 25.00 $\pm$ 8.39 21.72
+Layer-wise Caprese 84.80 29.17 $\pm$ 3.19 64.38 $\pm$ 7.18 27.50 $\pm$ 4.19 27.78
+E2E Caprese 85.40 30.00 $\pm$ 6.09 71.88 $\pm$ 5.54 29.17 $\pm$ 5.69 23.74
+CATS 89.20 34.17 $\pm$ 7.88 62.50 $\pm$ 2.04 37.50 $\pm$ 7.39 32.83
+Layer-wise Caprese 87.00 35.00 $\pm$ 4.30 70.00 $\pm$ 10.21 39.17 $\pm$ 3.19 38.38
+E2E Caprese 90.00 33.33 $\pm$ 4.71 78.13 $\pm$ 5.15 47.50 $\pm$ 5.00 35.86
+DS-Qwen 14B 92.80 63.33 $\pm$ 2.72 90.63 $\pm$ 2.39 60.83 $\pm$ 4.19 52.53
+GRIFFIN 89.80 32.50 $\pm$ 7.39 80.63 $\pm$ 3.75 33.33 $\pm$ 3.85 41.92
+Layer-wise Caprese 90.80 41.67 $\pm$ 5.77 86.88 $\pm$ 4.27 43.33 $\pm$ 4.71 55.05
+E2E Caprese 89.20 39.17 $\pm$ 9.18 84.38 $\pm$ 6.57 40.00 $\pm$ 2.72 43.43
+CATS 92.80 58.34 $\pm$ 1.92 88.75 $\pm$ 1.44 55.00 $\pm$ 3.33 44.95
+Layer-wise Caprese 91.00 61.67 $\pm$ 5.77 88.75 $\pm$ 1.44 53.00 $\pm$ 7.20 50.51
+E2E Caprese 92.00 58.34 $\pm$ 9.62 87.50 $\pm$ 2.04 49.17 $\pm$ 4.19 51.01
 4 Experiments
-We showcase the effectiveness of Caprese at recovering much, if not all, of the math performance lost from efficient inference algorithms without sacrificing efficiency or performance on language tasks.
-In the following section (Section 4.1), we observe better math performance for instruct LLMs without losing their generalizability.
-Then, we study Caprese’s benefit on two axes of test-time scaling: number of generations (Section 4.2) and generation length with thinking models (Section 4.3).
+We showcase the effectiveness of Caprese at recovering much, if not all, of the reasoning performance lost from efficient inference algorithms without sacrificing efficiency or performance on language tasks.
+Beginning with sequence length scaling in Section 4.1, Caprese is able to significantly or completely recover lost reasoning performance from existing efficient algorithms in thinking LLMs.
+Next, we demonstrate Caprese’s benefit when scaling the number of generations, another axis of tes-time scaling, in Section 4.2.
+Then in Section 4.3, we investigate typical instruct LLM settings without test-time scaling where we again observe better math performance without losing their generalizability.
 Finally, we highlight Caprese’s latency and length improvements in Section 4.4.
-When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying sparse method and similarly for GRIFFIN.
+Although we focus on coupling Caprese with sparse FF methods, we show our method’s merit carries over to quantization in Appendix D, broadening the potential applications of Caprese.
+When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying compression method and similarly for GRIFFIN.
 Otherwise, we use “Caprese” for brevity.
-Unless specified, the FF intermediate feature sparsity is set at 50%.
-NOTE: For a more meaningful baseline, we only apply GRIFFIN to the first half of the model for all experiments since we observed much steeper drops in math performance if used for all layers.
+Unless specified, FF intermediate feature sparsity is set at 50%.
+For a more meaningful baseline, we only apply GRIFFIN to the first half of models as we observe steeper drops in math performance if used for all layers.
 CATS is applied to all layers.
-4.1 Instruct Models
-We first look into different instruct LLMs which are performative on non-reasoning tasks and have decent reasoning performance.
-We show Caprese is able to preserve math performance without sacrificing quality on pure language tasks like question answering.
-We test on the Llama 3 family [DJP+24] and Gemma 2 family [TRP+24] of models on 0-shot GSM8K [CKB+21], MATH [HBK+21], CoQA [RCM19], and QASPER [DLB+21], using CoT prompts for math tasks.
-Table 1 shows that Caprese is able to preserve most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math.
-In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese is able to effectively recover the lost performance.
-E2E Caprese is the most performative in the majority of math scenarios, suggesting a benefit beyond minimizing local error.
-Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A).
-The best performers for CoQA and QASPER are a toss-up, but all methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases).
-CATS and Caprese considerably improve GSM8K performance of Llama 3.1 8B Instruct, which has trouble following the prompt template with zero-shot.
-4.2 Scaling Best-of-$N$: Coverage
-Figure 3: Coverage and standard deviation of 140 samples from MATH as the number of generation attempts, NN, scales.
-We define $\text{Relative Compute Units}=N\times A$ where $A$ is the fraction of total parameters activated per input.
-$F$ sparsity set to 50%.
-Best viewed zoomed.
-Now, we see the generalizability of Caprese on the first axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$.
-Increasing the number of generations per prompt, $N$, is one way to scale inference compute for improved performance, as the probability of generating a correct answer increases.
-To evaluate, we find the coverage (Pass@$K$) on 140 samples from MATH.
-We use an oracle verifier to accurately assess the quality of the pool of generated responses.
-To combat the high variance when $K$ is close to $N$, we calculate the average coverage for $K=1,\dots,100$ across 10 independent pools of 100 generations for each sample.
-Factoring in the saved compute, E2E Caprese is able to have similar or better coverage scaling compared to the original model, as shown in Figure 3.
-Notably, E2E Caprese (GRIFFIN) on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
 4.1 Scaling Length: Thinking Models
-Model MATH-500 AIME 2024 AMC 2023 BRUMO 2025
-DS-Qwen 1.5B 79.40 30.00 $\pm$ 2.72 60.83 $\pm$ 7.47 25.83 $\pm$ 3.19
-GRIFFIN 42.00 2.50 $\pm$ 1.67 16.88 $\pm$ 1.25 1.67 $\pm$ 1.92
-Layer-wise Caprese 47.20 2.50 $\pm$ 1.67 28.13 $\pm$ 2.39 0.00 $\pm$ 0.00
-E2E Caprese 60.40 6.67 $\pm$ 2.72 34.38 $\pm$ 6.57 4.17 $\pm$ 1.67
-CATS 72.00 11.67 $\pm$ 6.94 37.50 $\pm$ 3.54 10.83 $\pm$ 4.19
-Layer-wise Caprese 73.80 15.83 $\pm$ 3.19 48.13 $\pm$ 3.75 13.33 $\pm$ 2.72
-E2E Caprese 74.80 20.83 $\pm$ 1.67 47.50 $\pm$ 3.54 20.00 $\pm$ 2.72
-DS-Qwen 7B 90.20 47.50 $\pm$ 7.87 75.00 $\pm$ 4.08 45.83 $\pm$ 4.19
-GRIFFIN 80.60 21.67 $\pm$ 4.30 62.50 $\pm$ 8.90 25.00 $\pm$ 8.39
-Layer-wise Caprese 84.80 29.17 $\pm$ 3.19 64.38 $\pm$ 7.18 27.50 $\pm$ 4.19
-E2E Caprese 85.40 30.00 $\pm$ 6.09 71.88 $\pm$ 5.54 29.17 $\pm$ 5.69
-CATS 89.20 34.17 $\pm$ 7.88 62.50 $\pm$ 2.04 37.50 $\pm$ 7.39
-Layer-wise Caprese 87.00 35.00 $\pm$ 4.30 70.00 $\pm$ 10.21 39.17 $\pm$ 3.19
-E2E Caprese 90.00 33.33 $\pm$ 4.71 78.13 $\pm$ 5.15 47.50 $\pm$ 5.00
-DS-Qwen 14B 92.80 63.33 $\pm$ 2.72 90.63 $\pm$ 2.39 60.83 $\pm$ 4.19
-GRIFFIN 89.80 32.50 $\pm$ 7.39 80.63 $\pm$ 3.75 33.33 $\pm$ 3.85
-Layer-wise Caprese 90.80 41.67 $\pm$ 5.77 86.88 $\pm$ 4.27 43.33 $\pm$ 4.71
-E2E Caprese 89.20 39.17 $\pm$ 9.18 84.38 $\pm$ 6.57 40.00 $\pm$ 2.72
-CATS 92.80 58.34 $\pm$ 1.92 88.75 $\pm$ 1.44 55.00 $\pm$ 3.33
-Layer-wise Caprese 91.00 61.67 $\pm$ 5.77 88.75 $\pm$ 1.44 53.00 $\pm$ 7.20
-E2E Caprese 92.00 58.34 $\pm$ 9.62 87.50 $\pm$ 2.04 49.17 $\pm$ 4.19
-Table 2: Thinking models’ 0-shot accuracies on math reasoning tasks.
-Sparsity is set at 50%, and $r=256$.
-AIME 2024 and AMC 2023 columns also include the sample standard deviation across 4 runs.
-Further improvements with reselection are shown in Table 3.
-Thinking models provide another axis of test-time scaling by augmenting the CoT length before giving a final answer.
-Because this entails very long generation lengths, it is critical that error accumulation across tokens be minimized.
-We test on DeepSeek-R1-Distill-Qwen (which we abbreviate to DS-Qwen) models [YYZ+24, GYZ+25] using the prompt templates and configurations from Open R1 [Fac25] (temperature and top-$p$ set to 0.6 and 0.95, respectively).
-For these models, we test 0-shot performance on MATH-500 [HBK+21, LKB+23], AIME 2024 [AIM25], AMC 2023 [AMC23], and BRUMO 2025 [BRU25] for a max generation length of 32768.
+Thinking models scale inference by augmenting the CoT length before giving a final answer.
+Since this entails long generation lengths, it is critical that error accumulation across tokens be minimized.
+We test on DeepSeek-R1-Distill-Qwen (which we abbreviate to DS-Qwen) models [57, 23] using Open R1 [20] prompt templates and configurations (temperature and top-$p$ set to 0.6 and 0.95, respectively).
+We test 0-shot performance on math and science tasks MATH-500 [24, 34], AIME 2024 [1], AMC 2023 [2], BRUMO 2025 [6], and GPQA [44], for a max generation length of 32768.
 Due to the small size of AIME 2024, AMC 2023, and BRUMO 2025, we evaluate them 4 times, reporting the average accuracy and sample standard deviation.
+We observe none of the methods have any significant effect on standard deviation in Table 2.
 From Table 1, Caprese is the most performative in most cases.
@@
 DS-Qwen 7B with CATS brings AMC 2023 accuracy down to 62.50% from 75.00%, but Caprese lifts performance above the original model’s to 78.13%, and similarly with DS-Qwen 7B with CATS and BRUMO 2025.
-All methods are more robust as the model size increases.
-For instance, CATS and GRIFFIN had very little impact on DS-Qwen 14B’s accuracy on MATH-500, with Caprese performing similarly.
-AIME 2024 and BRUMO 2025 challenge CATS and GRIFFIN with several degradation and partial recovery using our method.
+All methods are more robust as model size increases.
+AIME 2024 and BRUMO 2025 challenge CATS and GRIFFIN with severe degradation and partial recovery with Caprese.
 In the next section, we show a simple way to push this recovery even further with reselection.
+Table 2: Average thinking model performance standard deviation on AIME 2024, AIME 2023, and BRUMO 2025 from Table 1. There is no clear consistent relationship between Caprese and accuracy variance.
+DS-Qwen 1.5B DS-Qwen 7B DS-Qwen 14B Mean
+Full Model 4.46 5.38 3.10 4.31
+GRIFFIN 1.61 7.20 5.00 4.60
+Layer-wise Caprese 1.35 4.85 4.92 3.71
+E2E Caprese 3.65 5.77 6.16 5.19
+CATS 4.89 5.77 2.23 4.30
+Layer-wise Caprese 3.22 5.90 4.80 3.97
+E2E Caprese 2.64 4.95 5.28 4.29
 4.1.1 Enhanced Performance with Reselection
@@
 As these tokens propagate through each layer, we find the selection metric for the generated tokens $\bm{s}_{G}$ and update corresponding KV pairs.
-This is like verification in speculative decoding [CBI+23, LKM23].
+This is like verification in speculative decoding [7, 32].
 As $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use (as described in Section 2).
 This updates the pruned layers yet avoids prefill for all tokens.
-Table 3 shows a clear benefit of reselection (even if infrequent compared to speculative decoding) in Caprese by pushing the performance much closer to the full model’s as we decrease the number of decode steps between reselection rounds.
-Although GRIFFIN is generally more harmful to accuracy than CATS due to its structured pruning, paired with Caprese and reselection, Caprese (GRIFFIN) can exceed the performance of Caprese (CATS) and reach the full model performance.
+Table 3 shows a clear benefit of reselection (even if infrequent compared to speculative decoding) in Caprese by pushing the performance much closer to the full model’s as we decrease steps between reselection rounds.
+Although GRIFFIN is generally more harmful to accuracy than CATS due to its structured pruning, with reselection, Caprese (GRIFFIN) can exceed the performance of Caprese (CATS) and reach the full model performance.
+Table 3: E2E Caprese AMC 2023 accuracies and standard deviations when recalculating GRIFFIN pruning metrics and reselecting pruned neurons every $\rho$ decode steps.
+No reselection and the full model are special cases where $\rho=\infty$ and $\rho=0$, respectively.
 Model No Reselect $\rho=1024$ $\rho=256$ $\rho=64$ Full
@@
 DS-Qwen 14B 84.38 88.13 88.75 91.88 90.63
-Table 3: E2E Caprese AMC 2023 accuracies and standard deviations when recalculating GRIFFIN pruning metrics and reselecting pruned neurons every $\rho$ decode steps.
-No reselection and the full model are special cases where $\rho=\infty$ and $\rho=0$, respectively.
+Figure 3: Coverage and standard deviation of 140 samples from MATH as the number of generation attempts, NN, scales.
+We define $\text{Relative Compute Units}=N\times A$ where $A$ is the fraction of total parameters activated per input.
+$F$ sparsity set to 50%.
+Best viewed zoomed.
 Reselection is also possible with CATS but requires recomputing prefill for all tokens.
-The parameter to update is the hard thresholding parameter $\tau$, but because this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory.
+The parameter to update is the hard thresholding parameter $\tau$, but since this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory.
 In turn, prefill will need to be redone anytime we want to update $\tau$.
@@
 Given this and the lack of computational incentive, we focus primarily on reselection for GRIFFIN.
+4.2 Scaling Best-of-$N$: Coverage
+Now, we see the generalizability of Caprese on the second axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$.
+Increasing the number of generations per prompt, $N$, increases the probability of generating a correct answer, even for non-thinking models (e.g., instruct models).
+To evaluate, we find the coverage (Pass@$K$) on 140 samples from MATH.
+We use an oracle verifier to accurately assess the quality of the pool of generated responses.
+To combat the high variance when $K$ is close to $N$, we calculate the average coverage for $K=1,\dots,100$ across 10 independent pools of 100 generations for each sample.
+Factoring in the saved compute, E2E Caprese is able to have similar or better coverage scaling compared to the original model, as shown in Figure 3.
+Notably, E2E Caprese (GRIFFIN) on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
+Table 4: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA, QASPER, XSum, and CNN/DailyMail). $F$ sparsity is set to 50%.
+Model GSM8K MATH CoQA QASPER XSum CNN/DailyMail
+Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43 21.65 25.60
+GRIFFIN 7.13 5.42 56.05 14.11 21.13 25.47
+Layer-wise Caprese 13.72 6.62 56.07 13.40 20.65 26.18
+E2E Caprese 21.00 8.44 56.55 13.88 20.71 26.18
+CATS 19.18 7.54 54.40 13.88 21.10 24.73
+Layer-wise Caprese 18.65 8.28 55.58 14.35 20.94 25.35
+E2E Caprese 20.39 9.04 56.12 13.75 20.40 25.56
+Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45 23.22 26.20
+GRIFFIN 28.96 10.98 64.52 12.52 22.09 25.49
+Layer-wise Caprese 40.18 13.70 64.33 11.60 21.56 25.90
+E2E Caprese 44.66 16.96 64.83 12.35 21.26 26.04
+CATS 41.24 12.04 58.87 11.36 22.23 25.40
+Layer-wise Caprese 45.49 13.62 60.53 12.26 22.50 25.90
+E2E Caprese 46.85 14.40 61.72 12.36 21.58 25.40
+Llama 3.1 8B Instruct 53.98 13.94 63.88 15.16 21.97 25.98
+GRIFFIN 20.47 6.16 63.37 12.63 21.53 25.89
+Layer-wise Caprese 37.60 9.72 65.05 14.21 21.92 26.31
+E2E Caprese 51.40 12.64 65.50 15.35 22.05 26.42
+CATS 50.95 11.80 58.85 12.26 21.43 25.67
+Layer-wise Caprese 51.93 13.66 63.92 14.34 22.58 25.79
+E2E Caprese 58.00 13.88 64.27 14.42 22.08 26.13
+Gemma 2 2B Instruct 51.02 16.06 63.77 10.96 22.17 26.01
+GRIFFIN 33.74 11.32 63.28 11.07 18.27 22.24
+Layer-wise Caprese 42.53 12.32 63.77 10.75 21.63 26.37
+E2E Caprese 48.14 13.70 63.37 11.05 22.48 27.16
+CATS 34.42 10.56 61.53 10.11 21.97 26.46
+Layer-wise Caprese 46.32 13.90 61.92 10.82 22.10 26.23
+E2E Caprese 46.17 14.16 63.92 11.03 22.15 26.52
+Gemma 2 9B Instruct 78.17 27.64 63.78 9.91 23.98 26.46
+GRIFFIN 59.21 25.22 63.82 10.14 24.66 26.77
+Layer-wise Caprese 76.72 25.84 64.20 9.82 24.88 26.88
+E2E Caprese 76.65 25.30 64.42 9.92 24.89 25.47
+CATS 76.50 27.32 64.37 9.78 24.74 26.64
+Layer-wise Caprese 77.18 28.16 64.52 10.46 24.54 26.45
+E2E Caprese 77.18 28.00 64.87 10.02 24.65 26.33
+4.3 Instruct Models
+Next, we look into the performance of Caprese in the absence of inference scaling methods on instruct LLMs whcih are primarily tailored for language tasks.
+We show Caprese is able to preserve math performance without sacrificing quality on language tasks like question answering.
+We test Llama 3 [19] and Gemma 2 [51] models on 0-shot GSM8K [8], MATH [24], CoQA [43], QASPER [10], XSum [40], and CNN/DailyMail [25, 47].
+We use CoT prompts for math tasks.
+Table 4 shows Caprese preserves most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math.
+In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese effectively recovers the lost performance.
+E2E Caprese is the most performative in the majority of math scenarios.
+Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A).
+All methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases).
 4.4 Efficiency
-Caprese reduces generation latency.
-Table 4 shows the latencies of different generation settings.
+Caprese reduces generation latency (Table 5).
 Caprese cuts total latency and time to next token by >16% for the Qwen 2.5 14B architecture.
@@
 Metrics were collected on an NVIDIA L40 GPU using BF16 precision.
+Table 5: End-to-end generation latency (s) and average time to next token (ms) for Qwen 2.5 14B.
+For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively.
+As before, GRIFFIN is applied to the first half of the model, sparsity is 50%, and $r=256$.
 Setup End-to-end Latency (s) Avg. Time to Next Token (ms)
@@
 2048+8192 344.9 287.7 312.5 288.8 42 35 38 35
-Table 4: End-to-end generation latency (s) and average time to next token (ms) for GRIFFIN, LoRA, and Caprese using Qwen 2.5 14B.
-For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively.
-As before, GRIFFIN is applied to the first half of the model, sparsity is 50%, and $r=256$.
 4.4.1 Bonus: Effect on Natural Response Length
@@
 Caprese implicitly encourages brevity, often even producing shorter responses than from the full thinking model.
-Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference (except for cutting off generation at 32K tokens).
-Shown in Figure 4 with MATH-500, the shortest mean response length for all problem difficulties is either the full model or Caprese.
-Meanwhile, CATS consistently outputs the longest responses, averaging roughly 1K more across every sample.
-It is also interesting to note that with increasing problem difficulty, all models and methods naturally allocate more inference tokens towards answering the question.
+Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference.
+Shown in Figure 4, the shortest mean response length for all MATH-500 problem difficulties is either the full model or Caprese.
+CATS consistently outputs the longest responses, averaging $\sim$1K more across every sample.
+With increasing difficulty, all models and methods naturally allocate more inference tokens towards the output.
 Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DS-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit.
 This brevity of Caprese is more pronounced in larger models, proportionally, from a 5.8% reduction in tokens compared to the full DS-Qwen 7B model to an 8.5% reduction compared to the full DS-Qwen 14B model.
-The cause of this phenomenon is of interest for future work.
-See Appendix D for similar observations with GRIFFIN.
+See Appendix E for observations with GRIFFIN.
 5 Conclusion
-To combat the inefficiency of the long and brittle generation process associated with math reasoning, we introduce Caprese, a highly performant and efficient method with strong math and language capabilities that is compatible with a broad class of efficient FF algorithms.
-In the future, it would be interesting to evaluate its benefit in other reasoning domains besides math, explore the use of input-dependent low-rank layers, and investigate the knowledge contained in these low-rank layers.
+To combat the inefficiency of the long and brittle generation process associated with reasoning, we introduce Caprese, a performant and efficient method with strong reasoning and language capabilities that is compatible with a broad class of efficient FF algorithms.
+In the future, it would be interesting to evaluate its benefit in more reasoning domains, explore input-dependent low-rank layers, and investigate the knowledge contained in these low-rank layers.
 Caprese, along with future developments, pushes towards the long-term goal of degradation-free efficient LLM inference.
@@
 Caprese achieves the highest accuracy in slightly more cases than LoRA, but as per our discussion on LoRA in Section 3.5 and efficiency analysis in Section 4.4, LoRA’s main benefit is efficient training, not efficient inference.
+Table 6: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA, QASPER, XSum, and CNN/DailyMail) with LoRA.
+$F$ sparsity is set to 50%, $r=256$, and LoRA ranks are set to match the number of parameters that Caprese adds.
+Model GSM8K MATH CoQA QASPER XSum CNN/DailyMail
+Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43 21.65 25.60
+GRIFFIN 7.13 5.42 56.05 14.11 21.13 25.47
+Layer-wise LoRA 8.11 5.60 56.00 14.65 21.08 25.51
+E2E LoRA 20.47 10.10 56.10 14.41 21.09 25.62
+Layer-wise Caprese 13.72 6.62 56.07 13.40 20.65 26.18
+E2E Caprese 21.00 8.44 56.55 13.88 20.71 26.18
+Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45 23.22 26.20
+GRIFFIN 28.96 10.98 64.52 12.52 22.09 25.49
+Layer-wise LoRA 26.46 11.28 63.88 12.55 22.12 25.96
+E2E LoRA 42.91 16.50 64.67 12.63 21.81 26.05
+Layer-wise Caprese 40.18 13.70 64.33 11.60 21.56 25.90
+E2E Caprese 44.66 16.96 64.83 12.35 21.26 26.04
+Gemma 2 2B Instruct 51.02 16.06 63.77 10.96 22.17 26.01
+GRIFFIN 33.74 11.32 63.28 11.07 18.27 22.24
+Layer-wise LoRA 39.58 14.66 62.98 10.54 21.58 26.64
+E2E LoRA 44.66 16.71 63.03 10.81 22.45 26.35
+Layer-wise Caprese 42.53 12.32 63.77 10.75 21.63 26.37
+E2E Caprese 48.14 13.70 63.37 11.05 22.48 27.16
+Appendix D Caprese with Quantization
+To demonstrate recovery generalizability beyond sparse FF methods as the backend, we also equip QLoRA [13] with Caprese.
+Table 7 shows that E2E Caprese is also able to nearly recover the performance degraded by quantization in both math and language tasks.
+Table 7: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA, QASPER) with QLoRA as the backend method. $F$ weights are quantized to 4 bits and $r=256$.
 Model GSM8K MATH CoQA QASPER
 Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43
-GRIFFIN 7.13 5.42 56.05 14.11
-Layer-wise LoRA 8.11 5.60 56.00 14.65
-E2E LoRA 20.47 10.10 56.10 14.41
-Layer-wise Caprese 13.72 6.62 56.07 13.40
-E2E Caprese 21.00 8.44 56.55 13.88
+QLoRA 12.05 6.56 48.70 12.37
+Layer-wise Caprese 15.01 6.88 49.60 12.65
+E2E Caprese 21.61 9.98 53.52 13.54
 Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45
-GRIFFIN 28.96 10.98 64.52 12.52
-Layer-wise LoRA 26.46 11.28 63.88 12.55
-E2E LoRA 42.91 16.50 64.67 12.63
-Layer-wise Caprese 40.18 13.70 64.33 11.60
-E2E Caprese 44.66 16.96 64.83 12.35
-Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
-GRIFFIN 33.74 11.32 63.28 11.07
-Layer-wise LoRA 39.58 14.66 62.98 10.54
-E2E LoRA 44.66 16.71 63.03 10.81
-Layer-wise Caprese 42.53 12.32 63.77 10.75
-E2E Caprese 48.14 13.70 63.37 11.05
-Table 5: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER) with LoRA.
-$F$ sparsity is set to 50%, $r=256$, and LoRA ranks are set to match the number of parameters that Caprese adds.
-Appendix D GRIFFIN Response Lengths
+QLoRA 45.41 11.98 62.22 12.87
+Layer-wise Caprese 45.56 11.78 64.25 12.41
+E2E Caprese 48.82 14.14 64.05 12.26
+Appendix E GRIFFIN Response Lengths
 Complementing Figure 4 for CATS, we show the natural response lengths to MATH-500 samples in Figure 7.
@@
 Figure 7: Average number of response tokens for MATH-500 queries with increasing problem difficulty. The global averages are indicated by the dashed lines. Sparsity is set at 50%.
-Appendix E Training Hyperparameters
+Appendix F Rank of Learned Parameters
+In Figure 8, we plot the relative singular values for each learned product $\bm{L}\bm{R}$ from (7) in Llama 3.2 3B Instruct.
+Although we set the inner rank to be 256, we see that some layers in Caprese(CATS) are still very low rank.
+In comparison, Caprese(GRIFFIN) layers are relatively high rank, likely to due to the fact that GRIFFIN performs highly structured FF sparsification.
+From this, we hypothesize that GRIFFIN may better utilize an increased Caprese inner dimension.
+Figure 8: Relative singular values of learned Llama 3.2 3B Instruct Caprese layers. Blue lines are individual layers; red lines are averaged across layers.
+Appendix G Training Hyperparameters
 Table 8 lists the hyperparameter settings for training Caprese layers. The E2E learning rates lie in the interval [4e-6, 2e-4], where larger models tend to learn better with smaller learning rates.
+Table 8: Caprese layer-wise and E2E training hyperparameters.
 Layer-wise End-to-end
@@
 Warmup 2% 2%
-Table 6: Caprese layer-wise and E2E training hyperparameters.